A Complete Bibliography of Publications in Nordic Journal of Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

06 May 2013 Version 1.10

Title word cross-reference

2 [76, 236, 225]. 3 [160, 44, 105]. B [23]. d [157, 141]. d/2 [157]. H [227]. K [259, 274]. K = N [259]. O(log log n) [26]. O(n^{2}) [51]. p [103]. π [3]. q [253]. N [54]. t [55].

-ary [55]. -calculus [3]. -categories [225]. -Claw [157]. -colorable [76]. -coloring [227]. -covering [227]. -D [44]. -Grams [253]. -Internal [274]. -Matching [105]. -packing [236]. -State [259]. -trees [23].

2.0 [251].

= [223].

ABEL [99]. Abstracting [233]. Access [137, 78, 38]. across [320]. Action [169, 241, 175]. Activities [251]. Activity [302]. Acyclic [238]. Agents [131]. Aggregate [297]. Aggregation [297]. Agile [316]. Aided [180]. ALFP [210]. Algebra [183, 174]. Algebraic [180, 211]. Algorithm [215, 112, 217, 236, 143, 146, 218, 85, 139, 187, 94, 182, 49, 91, 147, 23, 11, 31, 28, 26]. Algorithms [140, 74, 254, 56, 101, 158, 43, 294, 192, 129, 177, 10, 30, 304, 38]. Aliases [114]. Alignment [57]. All-Pairs [94]. amb [224]. AML [79]. Analysis [301, 100, 271, 242, 243, 294, 199, 303, 286, 167, 34]. Analyzing [234]. ancestors [24]. Animation [122]. API [291]. Application [205]. Applications [224, 229, 200]. Approach [243, 217, 291, 199, 313, 57, 234, 293]. Approximate [67, 194, 83, 255, 253, 18]. Approximating [254, 178]. Approximation [74, 157, 310, 158, 43, 143, 146, 192, 91, 17]. Approximation-based [310]. Approximations [75, 27]. Arbiter [263]. Arbitrary [67, 255]. Arc [65]. Architecture [96, 278]. Architecture-Driven [96]. Arrangements [125]. Arrays [261]. ary [55]. ASL [237]. Aspects [50]. Assessment [135]. Asynchronous [197]. ATM [138]. Attacks [285]. Attribute [79, 90, 81]. Attributed [149]. Attributes [290]. Auction [139]. Author [213, 235, 257]. Automata [208, 144, 199, 87, 86, 223]. Automata-Theoretic [199]. Automatic [279, 7]. Availability [234]. Average [67]. Axiomatizing [284].

B [241, 288]. Backtrack [62]. Bad [319]. Barriers [52]. Based [204, 242, 205, 291, 313, 315, 174, 165, 317, 119, 167, 310, 11, 261, 296, 212]. before [159]. Behaviour [212]. Behaviour-oriented [212]. Behavioural [265]. Bendix [59]. Beta [287]. Beta-Eta [287]. between [114]. Bidirectional [220]. Bigraphs [284]. Bin [189, 83]. Bin-Packing [83]. Binary [239]. Binding [284]. Bipartite [51]. Bisimulation [111, 225, 40]. Bit [218]. Bit-Vector [218]. Blinding [286]. Block [216, 306]. Blocks [185]. Blood [195]. Boppana [28]. Boppana-Halldórsson [28]. Bottleneck [192]. Bounded [64, 107, 163, 27, 18]. bounded-degree [27]. Bounding [88]. Bounds [305, 70, 127, 84, 300]. Bridge [114]. Broadcasting [61, 19]. Browsers [320]. Bucket [67]. Building [185, 114]. bundles [14]. Buses [47].

Cache [273]. Calculus [197, 308, 222, 37, 1, 20, 3]. Capturing [264]. Card [263]. Cardinality [51]. Cares [219]. Case [271, 241, 155, 274, 250, 138, 117]. Categorically [136]. categories [225]. Cell [125]. Cellular [309]. certifying [304]. Chain [298]. Change [123, 153]. Changing [88]. Characterization [306]. Checking [310, 173, 36, 21]. Chi [303]. Chip [278]. Chordal [63, 32]. chords [25]. Church [191]. Circular [65]. Circumvent [191]. Clairvoyance [128]. Class [272, 166, 279]. Classes [264]. Claw [157]. Clique [187, 9]. Clock [133]. Clones [319]. Closed [223]. closest [30, 31]. Closure [70]. Clustering [158, 13]. Code [208, 319, 291]. Coinserters [232]. colorable [76]. colored [31]. Coloring [76, 140, 227]. colors [76]. Coloured [270, 271, 272]. COM [117]. Combination [209]. Command [313]. Common [43, 49, 91, 24]. Communication [101]. Comonads [184]. Compact [86]. Comparing [138]. Comparison [276]. Comparisons [299, 300]. Competitive [189]. Compilation [121]. Compiler [80]. Compilers [248]. complete [17, 89]. Completion [59, 7]. Complexity [50, 145, 71, 307, 101, 106, 264, 300, 8, 17, 5]. Component [314]. Composition [36]. Compositional [265]. Comprehensibility [290]. Compression [50, 216]. Computational [132, 93, 5]. compute [26]. Computer [180]. Computers [47]. Computing [218, 97, 69, 246]. Concurrency [198, 38]. Concurrent [96, 200, 245, 39]. Conditional [121]. congruence [42]. Congruences [225]. Conic [142]. Connected [47]. Conscious [273]. Constant [64]. Constrained [73]. Constraint [112, 75, 200, 165, 167, 39]. Constraint-Based [165, 167]. Constraints [298]. Construction [256, 220, 23]. Constructive [98]. context [21]. context-free [21]. Continuous [169]. Control [197, 241, 171, 137, 296, 11].

Controlling [38]. Conventions [81]. Convex [160]. Coordination [131]. Corecursion [266]. Cores [134]. corridor [12]. Cost [242]. Cost-Based [242]. Countries [247, 115, 148]. Covering [106, 227]. Crochemore [217]. Cryptographic [111, 270]. CSP [209, 223]. CSP-OZ-DC [209]. Cut [145].

D [160, 44]. Dahm [195]. Damerau [218]. Data [50, 64, 209, 57, 273, 46, 10]. DC [209]. Decidability [197]. decomposition [36]. Deductive [282]. Degree [64, 27]. Delaunay [97]. Delegating [80]. Denotation [200]. Dependences [114]. Dependencies [78]. Dependent [231]. Depth [160]. Deques [229]. Derivation [180, 300]. Deriving [285, 225]. Described [144]. Descriptions [85]. Design [270, 271, 312, 294, 278, 73, 122]. Design/ CPN [270]. Detection [279]. Determinants [250]. Deterministic [260]. Development [118, 281, 116, 185, 205, 117, 247, 315, 115, 148, 316, 245, 288]. Diagrams [290, 302, 133, 48, 84, 52, 279]. diameter [9, 6]. Diameters [158]. Dictionaries [58, 155]. Difference [133, 298]. Differences [112]. different [305]. Dimensional [50, 144]. Dimensions [141, 176]. Direct [56]. Directed [238]. Directions [252]. Directives [57]. Disjoint [65, 14]. Distances [218]. Distinguishing [45]. Distributed [101, 57, 174, 177]. Distribution [129]. Distribution-Sensitive [129]. Distributive [112]. Document [239]. Documentation [151]. Dominance [56]. Dominating [63, 104]. domination [32, 8]. domination-type [8]. Don’t [219]. Down [159]. Driven [96, 293]. Dual [189]. Duration [308]. Dynamic [254, 256, 120, 202, 70, 137, 119, 11, 24].

Edge [156]. Edit [218]. Editor [154, 110, 214, 311, 280, 77, 16, 201, 318, 289, 130, 22, 162, 29]. Editors [196, 179, 207, 230, 35, 68, 275, 124, 247, 269, 115, 221, 148, 262, 168, 240]. Effectiveness [250]. Efficiency [322, 34]. Efficient [256, 112, 217, 294, 85, 155, 19, 31, 38]. Eiffel [321]. Element [299]. Elucidative [150, 203]. Embedded [118, 174, 296]. Embeddings [268]. Encoding [261]. Enforcing [81]. Engineering [276, 204, 80, 312, 132, 293]. Enhancing [316]. Enumerating [176]. Enumeration [125]. Environment [204, 116, 205, 247, 278, 115, 148, 119, 245, 82]. Episode [238]. epsilon [223]. epsilon-automata [223]. Equal [307]. equations [54]. Estimation [314]. Eta [287]. Euclidean [105]. Evaluation [295, 188]. Evolving [172]. exchange [10]. Executable [277]. Experiments [204, 139]. Explicit [290]. Exploration [193, 278]. Expression [255]. Expressive [222, 283]. Extensibility [4, 119]. Extensions [81]. Extreme [206, 176].

Fabric [138]. Facility [228]. Factor [192]. Factorization [86]. Family [51]. Fan [107]. Fan-Out [107]. Fanout [67]. Fast [47, 217, 143]. Fatal [59]. Fault [61, 38]. Fault-Tolerant [61, 38]. Featback [316]. Features [152]. Filter [299]. Finding [47, 25, 142, 9]. Finite [144, 86]. Finite-State [86]. First [166]. First-Class [166]. Fixpoint [112]. Flaw [285]. Fledged [321]. Flexible [49]. Flow [108, 14]. fly [244]. forbidden [304]. Foreword [196, 154, 110, 179, 207, 230, 214, 311, 280, 68, 77, 275, 124, 247, 269, 115, 221, 148, 201, 318, 262, 289, 130, 168, 240, 162, 35, 16, 22, 29]. fork [20]. Form [287]. Formal [281, 282, 132, 134, 303, 288]. Formalising [308]. Formalism [302]. Formulae [45]. Forward [301]. FPGA [134]. Fragmented [151]. Fragments [292]. Frameworks [276].

Free [157, 21]. Friendly [198]. Fully [321]. Fully-Fledged [321]. Functional [163]. Functions [90, 99]. Fusion [180]. Future [252].

Gain [88]. Games [40, 182]. Gaps [194]. General [170]. Generalized [227, 2, 104, 32]. Generate [291]. Generating [55]. Generation [137]. Generic [231, 243]. Genetic [317]. Geodesic [52]. Geometric [186, 297, 228, 13]. Geometry [93]. Giotto [296]. Good [319, 322]. Grammars [79, 90, 149]. Grams [253]. Graph [106, 238]. Graphs [140, 157, 156, 63, 104, 108, 65, 94, 27, 32, 28, 8]. greatest [54]. Guest [196, 154, 110, 179, 207, 230, 214, 311, 280, 35, 68, 275, 124, 247, 269, 115, 221, 148, 201, 318, 262, 289, 130, 168, 240, 22, 162]. Guided [171]. Guidelines [320].

HALDEN [237]. Halldórsson [28]. Hard [28, 18]. Hardness [273]. Hardware [303]. Hashing [100]. Height [95]. Heuristics [7, 306]. hierarchical [17]. Hierarchies [95]. Hierarchy [137]. Higher [99, 108, 176, 287, 15]. Higher-Order [108, 287]. history [211]. HOL [242, 3, 138]. Hybrid [169, 175, 246]. hypergraphs [76]. Hyperplane [142, 125]. Hypertextual [78].

Identifying [152]. Image [216]. Images [144]. Implementation [241, 216, 174, 46]. Implementing [248]. Improved [27, 193, 244, 127]. In-Place [163, 53]. Incomplete [279]. Increasing [290]. Incremental [172, 164]. Independence [198]. Independence-Friendly [198]. Independent [157, 274, 27]. Index [213, 235, 257]. Induced [147, 304]. induction [7]. Inductive [136]. Industrial [250]. Inference [89, 165]. Information [309]. Inheritance [120, 202, 321]. Input [189]. Instances [279]. Integer [255].

Integral [14]. Inter [167].

Inter-Procedural [167]. Interaction [224, 48]. Interactions [267]. Interactive [118]. Interdistances [254]. Interface [263, 265]. Interfaces [313]. Internal [274]. Internet [131]. Interpreters [185]. Intersection [63]. Interval [94]. Intervals [183]. Introducing [99]. Invariance [246]. Invariant [226]. Involving [297]. Isabelle [281]. Issues [252].

J2EE [205]. Java [248]. JavaScript [320]. Jobs [159]. Jumplists [260].

kernel [26]. Kernels [246]. Key [137].

Know [322]. Knuth [59].

label [305]. lambda [37, 1]. Language [118, 80, 283, 120, 174, 268]. Languages [264, 109]. layout [305]. Learning [90]. Least [54]. Length [261]. Levenshtein [218]. limits [37]. Line [189, 83, 93, 293]. Linear [100, 61, 304, 175, 147, 25, 2, 97]. Linear-time [304]. Lists [258]. Local [21]. Location [70]. Logarithmic [190]. Logic [198, 265, 113, 200]. logical [39]. logics [40]. Long [287]. Longest [49]. Loop [283]. Lossless [216]. Lower [70, 300]. lowest [24].

Machines [57]. Main [58]. Maintaining [71, 190, 252]. Maintenance [83]. Making [290]. Management [123, 153, 316, 293]. Manhattan [178]. Mapping [44]. Matching [215, 216, 51, 194, 188, 70, 105, 144, 85, 226, 87, 287, 253, 147]. Matchings [102]. Matrices [161]. MAXCLIQUE [28]. Maximal [183, 236]. Maximum [74, 157, 145, 299, 187, 147, 69]. Maximum-Minimum [299]. Maximum-Weight [187]. McCarthy [224]. MDA [317, 250]. MDG [138]. Mechanised [242]. mechanized [3]. Median [47, 258]. Median-Finding [47]. Memory [263, 57, 88, 58, 38]. Mendler [136].

Mendler-Style [136]. Mergesort [53]. Mesh [47]. Mesh-Connected [47]. Meshes [44, 33]. Message [92]. Metamodel [315]. Method [111, 58, 316, 10, 1]. Methods [67, 322]. minimization [18]. Minimize [158]. Minimum [71, 63, 72, 146, 178, 141, 299, 91, 9]. minimum-diameter [9]. Mismatches [215]. ML [79, 166]. Mobile [309]. Modal [198]. Mode [258]. Model [169, 310, 294, 313, 315, 317, 293, 292, 173, 36, 21]. Model-Based [313, 315]. Model-Checking [310]. Model-Driven [293]. Modeling [270, 276]. Modelled [272]. Modelling [132]. Models [290, 277, 315, 305, 33]. Modular [185]. Modularity [80]. Modules [88]. Moving [84]. Multi [226, 126]. Multi-Threaded [126]. Multi-Track [226]. Multicasting [107]. Multimessage [107]. Multiple [202, 320]. Multiprocessors [88].

Napier88 [82]. Navigation [229]. NC [140]. negative [42]. Neighborhoods [143]. Nets [301, 270, 271, 272, 224]. Network [178, 73]. Networks [309, 228, 305, 19, 10, 5]. neural [5]. No [299]. NoC [288]. Nodes [251]. non [6]. non-reversing [6]. Nondeterminism [267]. Nondeterministic [46]. noninterleaving [40]. Nordic [247, 115, 148]. Normal [287]. Note [215, 259, 217, 195, 15]. NP [89]. NP-complete [89]. Number [88, 76]. Numerals [191].

Oberon [4]. Object [271, 290, 152, 113, 122]. Object-Oriented [271, 152, 113, 122]. Objects [80, 1]. Obnoxious [228]. Off [155]. Offs [92, 6, 127]. old [211]. On-Line [189, 83]. On-the-fly [244]. Online [295, 126]. Open [153, 183]. Operational [60, 42]. Operations [297]. Optimal [161, 216, 56, 101, 95, 94, 23, 33]. Optimizing [282]. Order [99, 108, 287, 15].

Ordered [253]. Ordinary [142, 93]. Oriented [271, 152, 113, 122, 212]. Otway [285]. Output [125]. Output-Sensitive [125]. overlapped [34]. Overview [119]. OZ [209].

Packing [189, 83, 236]. Paging [126]. pair [30, 31]. Pairs [94]. Parallel [55, 216, 56, 294, 146, 94, 24, 167, 10, 38]. Parameterized [181]. Parametric [72, 102, 173]. Parentheses [70]. Parity [182]. Partial [190, 188]. Partially [64]. Partitioning [104]. Path [305, 139]. Paths [65, 94]. Pattern [188, 205, 144, 85, 291, 226, 317]. Pattern-Based [205, 291, 317]. Pattern-Matching [144, 85]. Patterns [249, 279]. CPN [270]. Penetration [160]. Permutation [140, 102]. Persistency [181]. Persistent [64, 82]. Petri [301, 270, 271, 272]. pi [197, 222]. pi-Calculus [197, 222]. Piles [229]. Place [163, 53]. Placement [273]. Placing [228]. Planar [70]. Point [70, 31]. Pointer [114]. Points [176, 84]. Polyadic [222]. polygon [25, 26]. Polygonal [246]. Polygons [193]. Polynomially [18]. Polytopes [160]. Power [222, 283, 103]. Power- [103]. Powerful [128]. Practical [277, 53]. PRAM [33]. predicates [42]. premises [42]. Presence [62, 52]. Presentation [122, 151]. Preserve [268]. Priority [229]. PRO [294]. Probabilistic [242, 41, 135]. Problem [197, 74, 145, 103, 92, 105, 139, 75, 93, 192, 187, 49, 91, 31]. Problems [186, 51, 56, 72, 297, 144, 95, 106, 73, 30, 12, 18, 17, 13, 8]. Procedural [152, 167]. Process [183, 315, 174, 250]. Process-Algebra [174]. Processes [209, 36, 21, 41]. processors [10]. Product [293]. Product-Line [293]. Profiles [249]. Program [180, 149, 78, 114, 203, 252]. Programming [256, 81, 264, 247, 174, 115, 200, 150, 148, 119, 82, 298, 39, 206].

Programs [231, 242, 171, 181, 113, 167]. Progress [183]. Proofs [231, 7]. Propagating [112]. Properties [224, 234]. Proportion [66]. Protocol [285]. Protocols [111, 270, 272, 234]. PSPACE [17]. PSPACE-complete [17]. pushdown [36]. Pyramids [44].

Quadtree [97]. quadtrees [34]. Quality [315]. quartic [69]. Queries [239, 295, 258]. Query [297]. Queues [229].

Radar [263]. Ragnarok [116]. Randomized [215, 71, 75, 182, 126, 30, 31, 28]. Range [297, 258]. Range-Aggregate [297]. Ratio [189]. Reachability [301, 197, 199]. Real [183, 243, 45, 314, 173]. Real-Time [183, 243, 314, 173]. Realization [250]. Reasoning [302]. Rebalancing [58]. recognition [304]. Recognizing [156]. Rectilinear [193, 52]. Recursion [170, 184]. Recursive [113]. Reducing [274]. Reduction [95]. Rees [285]. Refactoring [277]. Reference [149]. Refinement [135]. Refinements [233, 317]. Refining [267]. Reflective [313]. Regular [295, 310, 164, 255]. Regularities [219]. Relating [212]. Relational [78]. Reliability [135]. reliable [19]. removal [27]. Repetitions [217]. replication [11]. Representation [164]. Requirement [172]. ReRAGs [248]. Research [247, 115, 148]. Restricted [189]. Reusability [80]. Reverse [321]. reversing [6]. Revisited [100, 181, 188]. Rewriting [232, 285]. Rigourous [245]. Ring [259]. Roles [290]. Roundings [161]. Routing [102]. RPPS [199]. Run [261]. Run-Length [261].

Safe [202]. Safety [268]. Satisfaction [75]. Satisfiability [74]. Scalable [294]. Scenarios [291]. Schedulability [243]. Scheduling [159, 308]. Scheme [137].

Schemes [184]. Searching [62, 255]. Security [234]. Segmented [47]. Selection [127]. Self [259, 236]. Self-stabilization [259]. Self-stabilizing [236]. Selftypes [89]. Semantic [185, 90]. Semantics [60, 42]. Sensitive [129, 125]. Separability [156]. Separated [186]. Sequences [161, 189]. Series [186]. Service [312, 73]. Service-Constrained [73]. Set [157, 274]. Sets [186, 63, 104, 27, 31]. Shared [88]. Short [85]. Shortest [43, 139, 94]. Shut [159]. Shut-Down [159]. Similarity [186]. Simple [170, 30, 89]. Simplified [300]. SIMS [312]. simulation [33]. simulations [41]. Size [314]. Slicing [282]. Software [276, 204, 153, 116, 152, 132, 247, 314, 315, 115, 148]. Solution [83]. Solutions [287, 54]. Solvable [51]. Solver [210]. Some [13, 306, 211]. Sort [66]. Sorting [229, 306, 300, 2]. Source [153, 109]. Source-to-Source [109]. Space [256, 217, 163, 127, 322]. Space-Efficient [256, 217]. Spanning [72, 141, 274]. Sparse [19]. Sparsification [72]. specialization [1]. Specification [98, 209, 172, 181, 303]. Specifications [282, 265, 211, 46, 17]. Specifying [315]. Speed [128]. Split [66]. Splitting [173]. stabilization [259]. stabilizing [236]. Staging [268]. Standard [166]. State [259, 86, 322, 212]. State-based [212]. Static [286]. Statically [120]. Steiner [103]. Steps [59]. Stop [272]. Stop-and-Wait [272]. Storage [67]. Streaming [299]. Stretch [192]. String [215, 194, 219, 87, 253]. String-Matching [87]. Strings [188, 85, 226]. Strong [101]. Structural [279, 7]. Structure [274]. Structured [281, 95, 251, 46, 42]. Structures [64, 78, 166]. Structuring [177]. Study [271, 241, 250, 138]. Style [136]. sub [69]. sub-quartic [69]. Subexponential [182]. subgraph [27]. subgraphs [304]. sublinear [76]. Subsequence [49, 238]. Subsets [307]. subtrees [69]. Subtyping [98, 212]. Subworkshop [123, 153, 206]. Succinct [261, 210]. Suffix [220, 261]. Sum [307, 158]. Summary [123, 153, 206]. Sums [190]. Supersequence [43]. Supertree [91]. Support [204]. Supporting [315, 19]. survey [5]. Switch [138]. switching [305]. Symmetry [62]. Synchronisation [222]. Synthesis [208, 263, 171]. Synthesize [57]. Synthesizing [45]. System [163, 278, 165, 4]. Systematic [57]. Systems [169, 118, 241, 96, 112, 45, 172, 199, 303, 174, 296, 175, 245, 246, 288].

tabular [10]. Tasks [243]. Technique [81]. Techniques [209]. Technologies [131]. Temporal [200]. Termination [60]. Testers [244]. theorem [42]. Theoretic [199]. Theories [308]. Theory [231, 170, 3]. thoughts [211]. Threaded [97, 126]. Threads [151]. Three [299]. Time [186, 64, 190, 51, 101, 183, 243, 45, 209, 92, 314, 127, 173, 147, 25, 2, 304, 97, 69, 26]. Time-Message [92]. Time-Optimal [101]. Time-Series [186]. Time-Space [127]. Timed [301, 208, 223, 303]. Timing [309, 300]. Tolerant [61, 38]. Tool [117, 315]. Tools [312, 122, 252]. topological [2]. Trace [265]. Track [226]. Tracking [309]. Trade [6, 92, 155, 127]. Trade-Off [155]. Trade-Offs [92, 6, 127]. Transducers [86]. Transfinite [266]. Transformation [233]. Transitive [70]. Translation [121, 109]. Transposition [226]. Tree [295, 310, 72, 103, 95, 58, 274, 305]. Tree-Structured [95]. Tree-Type [58]. Trees [55, 239, 227, 220, 258, 141, 164, 173, 147, 23, 9]. Triangulated [156]. Triangulation [146, 97]. True [198]. TSP [143]. Turnip [195]. Tutorials [203]. Two [160, 50, 63, 144]. Two-Dimensional [50, 144]. Type [231, 170, 163, 58, 285, 89, 165, 8]. Typed [120]. Types [113, 60, 136, 46].

Ugly [319]. UML

[290, 277, 302, 267, 249, 251, 279].

Unambiguous [292]. Uncountable [37]. Undefinedness [237]. Underspecification [267, 237]. Understanding [252]. Unification [249]. Unique [164]. Unison [92]. Universes [231]. unrooted [69]. Unweighted [94]. Update [64, 163]. Updating [58]. UPPAAL [171]. Upper [127, 84]. Usage [291]. Users [309]. Using [301, 270, 241, 72, 171, 299, 149, 248, 173, 309, 225]. Utilization [67].

Validating [320]. Validation [312]. Value [108]. Variability [293]. Variants [256]. Variations [307]. Vector [218]. Verification [301, 270, 272, 282, 96, 263, 244, 134, 296, 138, 10]. Verifying [177]. Versatile [165]. Via [232, 27, 102]. view [11, 39]. view-based [11]. visible [25]. Visual [109]. Visualization [149]. Volume [213, 235, 257]. Voronoi [84, 52]. Voronoıdiagrams [15]. vs [237].

Wait [272]. Weak [92]. Weakly [156, 25]. Web [204, 320]. Web-Based [204]. Weight [157, 146, 187]. Weights [255]. Well [186]. Well-Separated [186]. Widest [12]. Widest-corridor [12]. Work [216, 33]. Work-Optimal [216, 33]. Workshop [82]. Worst [155]. Worst-Case [155].

References

Fisher:1994:LCO

[1] Kathleen Fisher, Furio Honsell, and John C. Mitchell. A lambda calculus of objects and method specialization. Nordic Journal of Computing, 1(1):3–37, Spring 1994. CO-DEN NJCOFR. ISSN 1236-6064.

                                             URL http://www.cs.helsinki.fi/
                                             njc/References/fisherhm1994:3.html.

Hagerup:1994:GTS

[2] Torben Hagerup and Martin Maas. Generalized topological sorting in linear time. Nordic Journal of Computing, 1(1):38–49, Spring 1994. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/hagerupm1994:38.
html.

Melham:1994:MTC

[3] T. F. Melham. A mechanized theory of the π-calculus in HOL. Nordic Journal of Computing, 1(1):50–76, Spring 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/melham1994:50.
html.

Mossenbock:1994:EOS

[4] Hanspeter Mössenböck. Extensibility in the Oberon system. Nordic Journal of Computing, 1(1):77–93, Spring 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/mossenbock1994:
77.html.

Orponen:1994:CCN

[5] Pekka Orponen. Computational complexity of neural networks: a survey. Nordic Journal of Computing, 1(1):94–110, Spring 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/orponen1994:94.
html.

Bodlaender:1994:TON

[6] Hans L. Bodlaender, Gerard Tel, and Nicola Santoro. Trade-offs in nonreversing diameter. Nordic Journal of Computing, 1(1):111–134, Spring 1994.

CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/bodlaenderts1994:
111.html.

Lysne:1994:HCA

[7] Olav Lysne. Heuristics for completion in automatic proofs by structural induction. Nordic Journal of Computing, 1(1):135–156, Spring 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/lysne1994:135.
html.

Telle:1994:CDT

[8] Jan Arne Telle. Complexity of domination-type problems in graphs. Nordic Journal of Computing, 1 (1):157–171, Spring 1994. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/telle1994:157.html.

Blair:1994:FMD

[9] Jean R. S. Blair and Barry W. Peyton. On finding minimum-diameter clique trees. Nordic Journal of Computing, 1(2):173–201, Summer 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/blairp1994:173.
html.

Coolsaet:1994:TMV

[10] K. Coolsaet, V. Fack, and H. De Meyer. A tabular method for verification of data exchange algorithms on networks of parallel processors. Nordic Journal of Computing, 1(2):202–213, Summer 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/coolsaetfm1994:
202.html.

Jajodia:1994:VBD

[11] S. Jajodia, R. Mukkamala, and K. V. S. Ramarao. A view-based dynamic replication control algorithm. Nordic Journal of Computing, 1(2):214–230, Summer 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/jajodiamr1994:
214.html.

Janardan:1994:WCP

[12] Ravi Janardan and Franco P. Preparata. Widest-corridor problems. Nordic Journal of Computing, 1(2):231–245, Summer 1994. CODEN NJCOFR. ISSN

1236-6064.  URL http://www.cs.
helsinki.fi/njc/References/janardanp1994:
231.html.

Pferschy:1994:SGC

[13] Ulrich Pferschy, Rüdiger Rudolf, and Gerhard J. Woeginger. Some geometric clustering problems. Nordic Journal of Computing, 1(2):246–263, Summer 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/pferschyrw1994:
246.html.

Jansen:1994:IFD

[14] Klaus Jansen. Integral flow with disjoint bundles. Nordic Journal of Computing, 1(2):264–267, Summer 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/jansen1994:264.
html.

Krussel:1994:NHO

[15] John W. Krussel and Barry F. Schaudt. A note on higher order Voronoĭdiagrams. Nordic Journal of Computing, 1 (2):268–272, Summer 1994. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/krussels1994:268.
html.

Lingas:1994:EF

[16] Andrzej Lingas. Editor’s foreword. Nordic Journal of Computing, 1(3):273– ??, Fall 1994. CODEN NJCOFR. ISSN 1236-6064.

Marathe:1994:CAP

[17] Madhav V. Marathe, Harry B. Hunt, III, and S. S. Ravi. The complexity of approximation PSPACE-complete problems for hierarchical specifications. Nordic Journal of Computing, 1(3):275–316, Fall 1994. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/marathehr1994:275.

html. Selected papers of the 20th International Colloquium on Automata, Languages and Programming (ICALP 93) (Lund, 1993).

Kann:1994:PBM

[18] Viggo Kann. Polynomially bounded minimization problems that are hard to approximate. Nordic Journal of Computing, 1(3):317–331, Fall 1994. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/kann1994:317.html.

Selected papers of the 20th International Colloquium on Automata, Languages and Programming (ICALP 93) (Lund, 1993).

Chlebus:1994:SNS

[19] Bogdan S. Chlebus, Krzysztof Diks, and Andrzej Pelc. Sparse networks supporting efficient reliable broadcasting. Nordic Journal of Computing, 1(3):332–345, Fall 1994. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/chlebusdp1994:332.

html. Selected papers of the 20th International Colloquium on Automata, Languages and Programming (ICALP 93) (Lund, 1993).

Havelund:1994:FC

[20] Klaus Havelund and Kim Guldstrand Larsen. The fork calculus. Nordic Journal of Computing, 1(3):346–363, Fall 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/havelundl1994:

346.html. Selected papers of the 20th International Colloquium on Automata, Languages and Programming (ICALP 93) (Lund, 1993).

Hungar:1994:LMC

[21] Hardi Hungar and Bernhard Steffen. Local model checking for context-free processes. Nordic Journal of Computing, 1(3):364–385, Fall 1994. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/hungars1994:364.

html. Selected papers of the 20th International Colloquium on Automata, Languages and Programming (ICALP 93) (Lund, 1993).

Skyum:1994:GEF

[22] Sven Skyum. Guest Editor’s foreword. Nordic Journal of Computing, 1(4):387– ??, Winter 1994. CODEN NJCOFR. ISSN 1236-6064.

Becker:1994:NAC

[23] Peter Becker. A new algorithm for the construction of optimal Btrees. Nordic Journal of Computing, 1(4):389–401, Winter 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/becker1994:389.

html. Selected papers of the 4th Scandinavian Workshop on Algorithm Theory (SWAT ’94) (Århus, 1994).

Schenk:1994:PDL

[24] Eric Schenk. Parallel dynamic lowest common ancestors. Nordic Journal of Computing, 1(4):402–432, Winter 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/schenk1994:402.

html. Selected papers of the 4th Scandinavian Workshop on Algorithm Theory (SWAT ’94) (Århus, 1994).

Das:1994:FAW

[25] Gautam Das, Paul J. Heffernan, and Giri Narasimhan. Finding all weakly visible chords of a polygon in linear time. Nordic Journal of Computing, 1(4):433–457, Winter 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/dashn1994:433.

html. Selected papers of the 4th Scandinavian Workshop on Algorithm Theory (SWAT ’94) (Århus, 1994).

Schuierer:1994:TAC

[26] Sven Schuierer. An O(log log n) time algorithm to compute the kernel of a polygon. Nordic Journal of Computing, 1(4):458–474, Winter 1994. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/schuierer1994:458.

html. Selected papers of the 4th Scandinavian Workshop on Algorithm Theory (SWAT ’94) (Århus, 1994).

Halldorsson:1994:IAI

[27] Magnús M. Halldórsson and Jaikumar Radhakrishnan. Improved approximations of independent sets in bounded-degree graphs via subgraph removal. Nordic Journal of Computing, 1(4):475–492, Winter 1994. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/halldorssonr1994:

475.html. Selected papers of the 4th Scandinavian Workshop on Algorithm Theory (SWAT ’94) (Århus, 1994).

Peinado:1994:HGR

[28] Marcus Peinado. Hard graphs for the randomized Boppana-Halldórsson algorithm for MAXCLIQUE. Nordic Journal of Computing, 1(4):493–515, Winter 1994. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/peinado1994:493.

html. Selected papers of the 4th Scandinavian Workshop on Algorithm Theory (SWAT ’94) (Århus, 1994).

Ukkonen:1995:EF

[29] Esko Ukkonen. Editor’s foreword. Nordic Journal of Computing, 2(1):1–??, Spring 1995. CODEN NJCOFR. ISSN 1236-6064.

Golin:1995:SRA

[30] Mordecai Golin, Rajeev Raman, Christian Schwarz, and Michiel Smid. Simple randomized algorithms for closest pair problems. Nordic Journal of Computing, 2(1):3–27, Spring 1995. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.

                                        fi/njc/References/golinrss1995:3.
                                        html.

Kamakoti:1995:ERA

[31] V. Kamakoti, Kamala Krithivasan, and C. Pandu Rangan. An efficient randomized algorithm for the closest pair problem on colored point sets. Nordic Journal of Computing, 2(1):28–40, Spring 1995. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/kamakotikr1995:
                                        28.html.

Kratochvil:1995:GDC

[32] Jan Kratochvíl, Paul D. Manuel, and Mirka Miller. Generalized domination in chordal graphs. Nordic Journal of Computing, 2(1):41–50, Spring 1995. CODEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/kratochvilmm1995:
                                        41.html.

Leppanen:1995:WOS

[33] Ville Leppänen and Martti Penttonen. Work-optimal simulation of PRAM models on meshes. Nordic Journal of Computing, 2(1):51–69, Spring 1995. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/leppanenp1995:
                                        51.html.

Vassilakopoulos:1995:EAO

[34] Michael Vassilakopoulos, Yannis Manolopoulos, and Brigitte Kröll. Efficiency analysis of overlapped quadtrees. Nordic Journal of Computing, 2(1):70–84, Spring 1995. CODEN NJCOFR. ISSN

                                        1236-6064.  URL http://www.cs.
                                        helsinki.fi/njc/References/vassilakopoulosmk1995:
                                        70.html.

Jonsson:1995:GEF

[35] B. Jonsson and Joachim Parrow. Guest Editors’ foreword. Nordic Journal of Computing, 2(2):87–??, Summer 1995. CODEN NJCOFR. ISSN 1236-6064.

Burkart:1995:CDM

[36] Olaf Burkart and Bernhard Steffen. Composition, decomposition and model checking of pushdown processes. Nordic Journal of Computing, 2(2):89–125, Summer 1995. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/burkarts1995:89.
html.

DiGianantonio:1995:ULL

[37] Pietro Di Gianantonio, Furio Honsell, and Gordon Plotkin. Uncountable limits and the lambda calculus. Nordic Journal of Computing, 2(2):126–145, Summer 1995. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/gianantoniohp1995:
126.html.

Kanellakis:1995:CMA

[38] Paris C. Kanellakis, Dimitrios Michailidis, and Alex Allister Shvartsman. Controlling memory access concurrency in efficient fault-tolerant parallel algorithms. Nordic Journal of Computing, 2(2):146–180, Summer 1995. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/kanellakisms1995:
146.html.

Mendler:1995:LVC

[39] Nax P. Mendler, Prakash Panangaden, P. J. Scott, and R. A. G. Seely.

A logical view of concurrent constraint programming. Nordic Journal of Computing, 2(2):181–220, Summer 1995. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/mendlerpss1995:
181.html.

Nielsen:1995:GLN

[40] Mogens Nielsen and Christian Clausen. Games and logics for a noninterleaving bisimulation. Nordic Journal of Computing, 2(2):221–249, Summer 1995. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/nielsenc1995:221.
html.

Segala:1995:PSP

[41] Roberto Segala and Nancy Lynch. Probabilistic simulations for probabilistic processes. Nordic Journal of Computing, 2(2):250–273, Summer 1995. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/segalal1995:250.
html.

Verhoef:1995:CTS

[42] Chris Verhoef. A congruence theorem for structured operational semantics with predicates and negative premises. Nordic Journal of Computing, 2(2):274–302, Summer 1995. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/verhoef1995:274.
html.

Fraser:1995:AAS

[43] Campbell B. Fraser and Robert W. Irving. Approximation algorithms for the shortest common supersequence. Nordic Journal of Computing, 2(3):303–325, Fall 1995. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/fraseri1995:303.
html.

Chung:1995:MPD

[44] Kuo Liang Chung and Yu Wei Chen. Mapping pyramids into 3-D meshes. Nordic Journal of Computing, 2(3):326–337, Fall 1995. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/chungc1995:326.
html.

Godskesen:1995:SDF

[45] Jens Chr. Godskesen and Kim G. Larsen. Synthesizing distinguishing formulae for real time systems. Nordic Journal of Computing, 2(3):338–357, Fall 1995. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/godskesenl1995:338.
html.

Walicki:1995:SSI

[46] Michał Walicki and Manfred Broy. Structured specifications and implementation of nondeterministic data types. Nordic Journal of Computing, 2(3):358–395, Fall 1995. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/walickib1995:358.
html.

Chung:1995:FMF

[47] Kuo-Liang L. Chung. Fast medianfinding on mesh-connected computers with segmented buses. Nordic Journal of Computing, 2(4):397–406, Winter 1995. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/chung1995:397.
html.

Parrow:1995:ID

[48] Joachim Parrow. Interaction diagrams. Nordic Journal of Computing, 2(4):407–443, Winter 1995. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/parrow1995:407.
html.

Rick:1995:NFA

[49] Claus Rick. A new flexible algorithm for the longest common subsequence problem. Nordic Journal of Computing, 2(4):444–461, Winter 1995. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/rick1995:444.html.

Bodlaender:1995:CAT

[50] Hans L. Bodlaender, Teofilo F. Gonzalez, and Ton Kloks. Complexity aspects of two-dimensional data compression. Nordic Journal of Computing, 2(4):462–495, Winter 1995. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/bodlaendergk1995:
462.html.

Clausen:1995:FBC

[51] Jens Clausen and Jakob Krarup. A family of bipartite cardinality matching problems solvable in O(n^{2}) time. Nordic Journal of Computing, 2(4):496–501, Winter 1995. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/clausenk1995:496.
html.

Tsin:1996:GVD

[52] Yung H. Tsin and Cao-An Wang. Geodesic Voronoi diagrams in the presence of rectilinear barriers. Nordic Journal of Computing, 3(1):1–26, Spring 1996. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/tsinw1996:1.html.

Katajainen:1996:PPM

[53] Jyrki Katajainen, Tomi Pasanen, and Jukka Teuhola. Practical in-place mergesort. Nordic Journal of Computing, 3(1):27–40, Spring 1996. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/katajainenpt1996:
27.html.

Seidl:1996:LGS

[54] Helmut Seidl. Least and greatest solutions of equations over N. Nordic Journal of Computing, 3(1):41–62, Spring 1996. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/seidl1996:41.
html.

Akl:1996:GAT

[55] Selim G. Akl and Ivan Stojmenovic. Generating t-ary trees in parallel. Nordic Journal of Computing, 3(1):63–71, Spring 1996. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/akls1996:63.html.

Datta:1996:OPA

[56] Amitava Datta, Anil Maheshwari, and Jörg-Rüdiger R. Sack. Optimal parallel algorithms for direct dominance problems. Nordic Journal of Computing, 3(1):72–88, Spring 1996. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/dattams1996:72.
html.

Lin:1996:SAS

[57] C.-Z. Lin, C.-C. Tseng, Y.-L. Chen, and T.-W. Kuo. A systematic approach to synthesize data alignment directives for distributed memory machines. Nordic Journal of Computing, 3(2):89–119, Summer 1996. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/lintck1996:89.
html.

Malmi:1996:NMU

[58] Lauri Malmi. A new method for updating and rebalancing tree-type main memory dictionaries. Nordic Journal of Computing, 3(2):111–130, Summer 1996. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/malmi1996:111.
html.

Linnestad:1996:FSK

[59] Henrik Linnestad. Fatal steps of Knuth– Bendix completion. Nordic Journal of Computing, 3(2):131–143, Summer 1996. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/linnestad1996:
131.html.

Nielson:1996:OST

[60] Flemming Nielson and Hanne Riis Nielson. Operational semantics of termination types. Nordic Journal of Computing, 3(2):144–187, Summer 1996. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/

njc/References/nielsonn1996:144.
html.

Diks:1996:FTL

[61] Krzysztof Diks and Andrzej Pelc. Fault-tolerant linear broadcasting. Nordic Journal of Computing, 3 (2):188–201, Summer 1996. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/diksp1996:188.html.

Brown:1996:BSP

[62] Cynthia A. Brown, Larry Finkelstein, and Paul Walton Purdom, Jr. Backtrack searching in the presence of symmetry. Nordic Journal of Computing, 3(3):203–219, Fall 1996. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/brownfp1996:203.
html.

Eades:1996:TMD

[63] Peter Eades, Mark Keil, Paul D. Manuel, and Mirka Miller. Two minimum dominating sets with minimum intersection in chordal graphs. Nordic Journal of Computing, 3(3):220– 237, Fall 1996. CODEN NJCOFR. ISSN 1236-6064. URL http://www.cs.

helsinki.fi/njc/References/eadeskmm1996:
220.html.

Brodal:1996:PPD

[64] Gerth Stølting Brodal. Partially persistent data structures of bounded degree with constant update time. Nordic Journal of Computing, 3(3):238–255, Fall 1996. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/brodal1996:238.
html.

Natarajan:1996:DPC

[65] Sridhar Natarajan and Alan P. Sprague. Disjoint paths in circular arc graphs. Nordic Journal of Computing, 3(3):256– 270, Fall 1996. CODEN NJCOFR. ISSN 1236-6064. URL http://www.cs.

                                        helsinki.fi/njc/References/natarajans1996:
                                        256.html.

Chen:1996:PSS

[66] J.-C. Chen. Proportion split sort. Nordic Journal of Computing, 3(3):271– 279, Fall 1996. CODEN NJCOFR. ISSN 1236-6064. URL http://www.cs.

                                        helsinki.fi/njc/References/chen1996:
                                        271.html.

Ang:1996:AAS

[67] C.-H. Ang and H. Samet. Approximate average storage utilization of bucket methods with arbitrary fanout. Nordic Journal of Computing, 3(3):280–291, Fall 1996. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/angs1996:280.html.

Karlsson:1996:GEF

[68] R. Karlsson and Andrzej Lingas. Guest Editors’ foreword. Nordic Journal of Computing, 3(4):293–294, Winter 1996. CODEN NJCOFR. ISSN 1236-6064.

Lam:1996:CUM

[69] T. W. Lam, W. K. Sung, and H. F. Ting. Computing unrooted maximum subtrees in sub-quartic time. Nordic Journal of Computing, 3(4):295–322, Winter 1996. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/lamst1996:295.

html. The 5th Scandinavian Workshop on Algorithm Theory (SWAT ’96) (Reykjavík, 1996).

Husfeldt:1996:LBD

[70] Thore Husfeldt, Theis Rauhe, and Søren Skyum. Lower bounds for dynamic transitive closure, planar point location, and parentheses matching. Nordic Journal of Computing, 3(4):323–336, Winter 1996. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/husfeldtrs1996:

323.html. The 5th Scandinavian Workshop on Algorithm Theory (SWAT ’96) (Reykjavík, 1996).

Brodal:1996:RCM

[71] Gerth Stølting Brodal, Shiva Chaudhuri, and Jaikumar Radhakrishnan. The randomized complexity of maintaining the minimum. Nordic Journal of Computing, 3(4):337–351, Winter 1996. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/brodalcr1996:337.

html. The 5th Scandinavian Workshop on Algorithm Theory (SWAT ’96) (Reykjavík, 1996).

Fernandez-Baca:1996:USP

[72] David Fernández-Baca, Giora Slutzki, and David Eppstein. Using sparsification for parametric minimum spanning tree problems. Nordic Journal of Computing, 3(4):352–366, Winter 1996. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/fernandez-bacase1996:

352.html. The 5th Scandinavian Workshop on Algorithm Theory (SWAT ’96) (Reykjavík, 1996).

Marathe:1996:SCN

[73] M. V. Marathe, R. Ravi, and R. Sundaram. Service-constrained network design problems. Nordic Journal of Computing, 3(4):367–387, Winter 1996. CODEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/marathers1996:367.
                                        html.

Asano:1996:AAM

[74] Takao Asano, Takao Ono, and Tomio Hirata. Approximation algorithms for the maximum satisfiability problem. Nordic Journal of Computing, 3(4):388–404, Winter 1996. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/asanooh1996:388.

html. The 5th Scandinavian Workshop on Algorithm Theory (SWAT ’96) (Reykjavík, 1996).

Lau:1996:RAC

[75] Hoong Chuin Lau and Osamu Watanabe. Randomized approximations of the constraint satisfaction problem. Nordic Journal of Computing, 3(4):405–424, Winter 1996. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/lauw1996:405.html.

The 5th Scandinavian Workshop on Algorithm Theory (SWAT ’96) (Reykjavík, 1996).

Alon:1996:CCH

[76] Noga Alon, Pierre Kelsen, Sanjeev Mahajan, and Hariharan Ramesh. Coloring 2-colorable hypergraphs with a sublinear number of colors. Nordic Journal of Computing, 3(4):425–439, Winter 1996. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/alonkmr1996:425.

html. The 5th Scandinavian Workshop on Algorithm Theory (SWAT ’96) (Reykjavík, 1996).

Koskimies:1997:EF

[77] K. Koskimies. Editor’s foreword. Nordic Journal of Computing, 4(1):1–2, Spring 1997. CODEN NJCOFR. ISSN 1236-6064.

Paakki:1997:RPD

[78] J. Paakki, J. Koskinen, and A. Salminen. From relational program dependencies to hypertextual access structures. Nordic Journal of Computing, 4(1):3–36, Spring 1997. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/paakkiks1997:3.
html.

Efremidis:1997:AAG

[79] Sofoklis G. Efremidis, Khalid A. Mughal, John H. Reppy, and Lars Søraas. AML: Attribute grammars in ML. Nordic Journal of Computing, 4(1):37–65, Spring 1997. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/efremidismrq1997:
37.html.

Bosch:1997:DCO

[80] Jan Bosch. Delegating compiler objects: Modularity and reusability in language engineering. Nordic Journal of Computing, 4(1):66–92, Spring 1997. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/bosch1997:66.html.

Hedin:1997:AET

[81] Goerel Hedin. Attribute extensions – a technique for enforcing programming conventions. Nordic Journal of Computing, 4(1):93–122, Spring 1997. CODEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/hedin1997:93.html.

Sjoberg:1997:PWP

[82] Dag I. K. Sjøberg, Ray Welland, Malcolm P. Atkinson, Paul Philbrow, Cathy Waite, and Stewart Macneill. The persistent workshop — a programming environment for Napier88. Nordic Journal of Computing, 4 (1):123–149, Spring 1997. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/sjobergwapwm1997:
                                        123.html.

Gambosi:1997:LMA

[83] Giorgio Gambosi, Alberto Postiglione, and Maurizio Talamo. On-line maintenance of an approximate bin-packing solution. Nordic Journal of Computing, 4(2):151–166, Summer 1997. CODEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/gambosipt1997:151.
                                        html.

Roos:1997:NUB

[84] Thomas Roos. New upper bounds on Voronoi diagrams of moving points. Nordic Journal of Computing, 4 (2):167–171, Summer 1997. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/roos1997:167.html.

Karpinski:1997:EPM

[85] Marek Karpinski, Wojciech Rytter, and Ayumi Shinohara. An efficient pattern-matching algorithm for strings with short descriptions. Nordic Journal of Computing, 4 (2):172–186, Summer 1997. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/karpinskirs1997:
172.html.

Roche:1997:CFF

[86] Emmanuel Roche. Compact factorization of finite-state transducers and finite-state automata. Nordic Journal of Computing, 4(2):187–216, Summer 1997. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/roche1997:187.
html.

Mohri:1997:SMA

[87] Mehryar Mohri. String-matching with automata. Nordic Journal of Computing, 4(2):217–231, Summer 1997. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/mohri1997:217.
html.

Lundberg:1997:BGC

[88] Lars Lundberg and Håkan Lennerstad. Bounding the gain of changing the number of memory modules in shared memory multiprocessors. Nordic Journal of Computing, 4(3):233–258, Fall 1997. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/lundbergl1997:
233.html.

Palsberg:1997:TIS

[89] Jens Palsberg and Trevor Jim. Type inference with simple selftypes is NPcomplete. Nordic Journal of Computing, 4(3):259–286, Fall 1997. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/palsbergj1997:259.
                                        html.

Gyimothy:1997:LSF

[90] Tibor Gyimóthy and Tamás Horváth. Learning semantic functions of attribute grammars. Nordic Journal of Computing, 4(3):287–302, Fall 1997. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/gyimothyh1997:287.
                                        html.

Yamaguchi:1997:AAM

[91] Atsuko Yamaguchi, Koji Nakano, and Satoru Miyano. An approximation algorithm for the minimum common supertree problem. Nordic Journal of Computing, 4(3):303–316, Fall 1997. CODEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/yamaguchinm1997:
                                        303.html.

Israeli:1997:TMT

[92] Amos Israeli, Evangelos Kranakis, Danny Krizanc, and Nicola Santoro. Time-message trade-offs for the weak unison problem. Nordic Journal of Computing, 4(4):317–329, Winter 1997. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/israelikks1997:
                                        317.html.

Mukhopadhyay:1997:OLP

[93] Asish Mukhopadhyay, Alok Agrawal, and Ravi Mohan Hosabettu. On the ordinary line problem in computational geometry. Nordic Journal of Computing, 4(4):330–341, Winter 1997. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/mukhopadhyayah1997:
330.html.

Pal:1997:OPA

[94] Madhumangal Pal and G. P. Bhattacharjee. An optimal parallel algorithm for all-pairs shortest paths on unweighted interval graphs. Nordic Journal of Computing, 4(4):342–356, Winter 1997. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/palb1997:342.
html.

Kaser:1997:OHR

[95] Owen Kaser. Optimal height reduction problems for tree-structured hierarchies. Nordic Journal of Computing, 4(4):357–379, Winter 1997. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/kaser1997:357.
html.

Erdogmus:1997:ADV

[96] Hakan Erdogmus. Architecturedriven verification of concurrent systems. Nordic Journal of Computing, 4(4):380–413, Winter 1997. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/erdogmus1997:380.
html.

Krznaric:1998:CTQ

[97] Drago Krznaric and Christos Levcopoulos. Computing a threaded quadtree from the Delaunay triangulation in linear time. Nordic Journal of Computing, 5(1):1–18, Spring 1998. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/krznaricl1998:
                                        1.html.

Dahl:1998:SCS

[98] Ole-Johan Dahl, Olaf Owe, and Tore J. Bastiansen. Subtyping and constructive specification. Nordic Journal of Computing, 5(1):19–49, Spring 1998. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/dahlob1998:19.
                                        html.

Kristoffersen:1998:IHO

[99] Bjørn Kristoffersen and Ole-Johan Dahl. On introducing higher order functions in ABEL. Nordic Journal of Computing, 5(1):50–69, Spring 1998. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/kristoffersend1998:
                                        50.html.

Baeza-Yates:1998:ALH

[100] Ricardo A. Baeza-Yates and Héctor Soza-Pollman. Analysis of linear hashing revisited. Nordic Journal of Computing, 5(1):70–85, Spring 1998. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/baeza-yatess1998:
                                        70.html.

Dobrev:1998:CCS

[101] Stefan Dobrev and Peter Ružička. On the communication complexity of strong time-optimal distributed algorithms. Nordic Journal of Computing, 5(2):87–104, Summer 1998. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/dobrevr1998:87.
html.

Hoyer:1998:PPR

[102] Peter Høyer and Kim S. Larsen.

Parametric permutation routing via matchings. Nordic Journal of Computing, 5(2):105–114, Summer 1998. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/hoyerl1998:105.
html.

Ganley:1998:PST

[103] Joseph L. Ganley and Jeffrey S. Salowe. The power-p Steiner tree problem. Nordic Journal of Computing, 5(2):115– 127, Summer 1998. CODEN NJCOFR. ISSN 1236-6064. URL http://www.cs.

helsinki.fi/njc/References/1998:
115.html.

Heggernes:1998:PGG

[104] Pinar Heggernes and Jan Arne Telle.

Partitioning graphs into generalized dominating sets. Nordic Journal of Computing, 5(2):128–142, Summer 1998. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/heggernest1998:
128.html.

Johnsson:1998:EMP

[105] Mika Johnsson, Gábor Magyar, and Olli Nevalainen. On the Euclidean 3-matching problem. Nordic Journal of Computing, 5(2):143–171, Summer 1998. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.

                                        fi/njc/References/johnssonmn1998:
                                        143.html.

Kratochvil:1998:CGC

[106] Jan Kratochvíl, Andrzej Proskurowski, and Jan Arne Telle. On the complexity of graph covering problems. Nordic Journal of Computing, 5(3):173–??, Fall 1998. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/kratochvilpt1998:
                                        173.html.

Gonzalez:1998:BFM

[107] Teofilo F. Gonzalez. Bounded fanout multimessage multicasting. Nordic Journal of Computing, 5(3):196–??, Fall 1998. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/gonzalez1998:
                                        196.html.

Mossin:1998:HOV

[108] Christian Mossin. Higher-order value flow graphs. Nordic Journal of Computing, 5(3):214–??, Fall 1998. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/mossin1998:214.
                                        html.

Paakki:1998:SST

[109] Jukka Paakki and Antti-Pekka Tuovinen. Source-to-source translation of visual languages. Nordic Journal of Computing, 5(3):235–??, Fall 1998. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/paakkit1998:235.
                                        html.

Hankin:1998:GEF

[110] Chris Hankin. Guest Editor’s foreword. Nordic Journal of Computing, 5(4):265– ??, Winter 1998. CODEN NJCOFR. ISSN 1236-6064.

Abadi:1998:BMC

[111] Martín Abadi and Andrew D. Gordon. A bisimulation method for cryptographic protocols. Nordic Journal of Computing, 5(4):267–??, Winter 1998. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/abadig1998:267.
html.

Fecht:1998:PDE

[112] Christian Fecht and Helmut Seidl.

Propagating differences: an efficient new fixpoint algorithm for distributive constraint systems. Nordic Journal of Computing, 5(4):304–??, Winter 1998. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/fechts1998:304.
html.

Leino:1998:ROT

[113] K. Rustan M. Leino. Recursive object types in a logic of objectoriented programs. Nordic Journal of Computing, 5(4):330–??, Winter 1998. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/leino1998:330.
html.

Ross:1998:BBB

[114] John L. Ross and Mooly Sagiv. Building a bridge between pointer aliases and program dependences. Nordic Journal of Computing, 5(4):361–??, Winter 1998. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/rosss1998:361.
html.

Mughal:1999:GEF

[115] Khalid A. Mughal and Andreas L. Opdahl. Guest Editors’ foreword: Programming and software development environment research in the Nordic countries. Nordic Journal of Computing, 6 (1):1–??, Spring 1999. CODEN NJ-COFR. ISSN 1236-6064.

Christensen:1999:RSD

[116] Henrik Bærbak Christensen. The Ragnarok software development environment. Nordic Journal of Computing, 6(1):4–??, Spring 1999. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/christensen1999:4.
                                        html.

Kjelstrom:1999:CTC

[117] Kåre Kjelstrøm and Peter Petersen. A CASE tool for COM development. Nordic Journal of Computing, 6(1):22–??, Spring 1999. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/kjelstromp1999:22.
                                        html.

Bjarnason:1999:ILD

[118] Elizabeth Bjarnason, Görel Hedin, and Klas Nilsson. Interactive language development for embedded systems. Nordic Journal of Computing, 6(1):36–??, Spring 1999. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/bjarnasonhn1999:
                                        36.html.

Ron:1999:ODP

[119] Henrik Røn. An overview of a dynamic programming environment based on extensibility. Nordic Journal of Computing, 6(1):55–??, Spring 1999. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/ron1999:55.html.

Ernst:1999:DIS

[120] Erik Ernst. Dynamic inheritance in a statically typed language. Nordic Journal of Computing, 6(1):72–??, Spring 1999. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/ernst1999:72.
html.

Harsu:1999:TCC

[121] Maarit Harsu. Translation of conditional compilation. Nordic Journal of Computing, 6(1):93–??, Spring 1999. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/harsu1999:93.html.

Normark:1999:APT

[122] Kurt Nørmark, Lars Iversen, and Per Madsen. Animation and presentation tools for object-oriented design. Nordic Journal of Computing, 6(1):110–??, Spring 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/normarkim1999:110.
html.

Bendix:1999:SSC

[123] Lars Bendix and Ulf Asklund. Summary of the subworkshop on change management. Nordic Journal of Computing, 6 (1):129–??, Spring 1999. CODEN NJ-COFR. ISSN 1236-6064.

Lagergren:1999:GEF

[124] Jens Lagergren and Peter Bro Miltersen.

Guest Editors’ foreword. Nordic Journal of Computing, 6(2):135–??, Summer 1999. CODEN NJCOFR. ISSN 1236-6064.

Sleumer:1999:OSC

[125] Nora H. Sleumer. Output-sensitive cell enumeration in hyperplane arrangements. Nordic Journal of Computing, 6(2):137–??, Summer 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/sleumer1999:137.
html.

Seiden:1999:ROM

[126] Steven S. Seiden. Randomized online multi-threaded paging. Nordic Journal of Computing, 6(2):148–??, Summer 1999. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/seiden1999:148.
html.

Raman:1999:IUB

[127] Venkatesh Raman and Sarnath Ramnath. Improved upper bounds for time-space trade-offs for selection. Nordic Journal of Computing, 6(2):162–??, Summer 1999. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/ramanr1999:162.
html.

Berman:1999:SMP

[128] Piotr Berman and Chris Coulston. Speed is more powerful than clairvoyance. Nordic Journal of Computing, 6(2):181–??, Summer 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/bermanc1999:181.
html.

Sen:1999:DSA

[129] Sandeep Sen and Neelima Gupta.

Distribution-sensitive algorithms. Nordic Journal of Computing, 6(2):194–??, Summer 1999. CODEN NJCOFR. ISSN 1236-6064. URL http://www.cs.

helsinki.fi/njc/References/seng:
194.html.

Sere:1999:GEF

[130] Kaisa Sere. Guest Editor’s foreword.

Nordic Journal of Computing, 6(3):213– ??, Fall 1999. CODEN NJCOFR. ISSN 1236-6064.

Ciancarini:1999:CTI

[131] Paolo Ciancarini, Andrea Omicini, and Franco Zambonelli. Coordination technologies for Internet agents. Nordic Journal of Computing, 6(3):215–??, Fall 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/ciancarinioz1999:
215.html.

Haveraaen:1999:FSE

[132] Magne Haveraaen, Helmer André Friis, and Tor Arne Johansen. Formal software engineering for computational modelling. Nordic Journal of Computing, 6(3):241–??, Fall 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/haveraaenfj1999:
241.html.

Larsen:1999:CDD

[133] Kim G. Larsen, Justin Pearson, Carsten Weise, and Wang Yi. Clock difference diagrams. Nordic Journal of Computing, 6(3):271–??, Fall 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/larsenpwy1999:271.
html.

Lillieroth:1999:FVF

[134] Carl Johan Lillieroth and Satnam Singh. Formal verification of FPGA cores. Nordic Journal of Computing, 6(3):299–??, Fall 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/lillieroths1999:
299.html.

Troubitsyna:1999:RAT

[135] Elena A. Troubitsyna. Reliability assessment through probabilistic refinement. Nordic Journal of Computing, 6(3):320–??, Fall 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/troubitsyna1999:
320.html.

Uustalu:1999:MSI

[136] Tarmo Uustalu and Varmo Vene. Mendler-style inductive types, categorically. Nordic Journal of Computing, 6(3):343–??, Fall 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/uustaluv1999:343.
html.

Hwang:1999:DKG

[137] Min-Shiang Hwang. A dynamic key generation scheme for access control in a hierarchy. Nordic Journal of Computing, 6(4):363–??, Winter 1999. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/hwang1999:363.
html.

Tahar:1999:CHM

[138] Sofiène Tahar and Paul Curzon. Comparing HOL and MDG: a case study on the verification of an ATM switch fabric. Nordic Journal of Computing, 6(4):372–??, Winter 1999. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/taharc1999:372.
html.

Larsen:1999:EAA

[139] Jesper Larsen and Ib Pedersen. Experiments with the auction algorithm for the shortest path problem. Nordic Journal of Computing, 6(4):403–??, Winter 1999. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/larsenp1999:403.
html.

Andreou:1999:NCA

[140] Maria Andreou and Stavros D. Nikolopoulos. NC coloring algorithms for permutation graphs. Nordic Journal of Computing, 6(4):422–??, Winter 1999. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/andreoun:422.html.

Krznaric:1999:MST

[141] Drago Krznaric, Christos Levcopoulos, and Bengt J. Nilsson. Minimum spanning trees in d dimensions. Nordic Journal of Computing, 6(4):446–??, Winter 1999. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/krznaricln:446.
html.

Devillers:1999:FOC

[142] Olivier Devillers and Asish Mukhopadhyay. Finding an ordinary conic and an ordinary hyperplane. Nordic Journal of Computing, 6(4):462–??, Winter 1999. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/devillersm:462.
html.

Gudmundsson:1999:FAA

[143] Joachim Gudmundsson and Christos Levcopoulos. A fast approximation algorithm for TSP with neighborhoods. Nordic Journal of Computing, 6(4):469–??, Winter 1999. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/gudmundssonl:469.
html.

Karhumaki:2000:PMP

[144] Juhani Karhumäki, Wojciech Plandowski, and Wojciech Rytter. Pattern-matching problems for two-dimensional images described by finite automata. Nordic Journal of Computing, 7(1):1–??, Spring 2000. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/karhumakipr2000:
1.html.

Bodlaender:2000:CMC

[145] Hans L. Bodlaender and Klaus Jansen. On the complexity of the maximum cut problem. Nordic Journal of Computing, 7(1):14–??, Spring 2000. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/bodlaenderj:14.
html.

Gudmundsson:2000:PAA

[146] Joachim Gudmundsson and Christos Levcopoulos. A parallel approximation algorithm for minimum weight triangulation. Nordic Journal of Computing, 7(1):32–??, Spring 2000. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/gudmundssonl:32.
html.

Zito:2000:LTM

[147] Michele Zito. Linear time maximum induced matching algorithm for trees. Nordic Journal of Computing, 7(1):58–??, Spring 2000. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/zito2000:58.html.

Opdahl:2000:GEF

[148] Andreas L. Opdahl and Dag I. K.

Sjøberg. Guest Editors’ foreword: Programming and software development environment research in the Nordic countries. Nordic Journal of Computing, 7 (2):65–??, Summer 2000. CODEN NJ-COFR. ISSN 1236-6064.

Magnusson:2000:PVU

[149] Eva Magnusson and Görel Hedin. Program visualization using reference attributed grammars. Nordic Journal of Computing, 7(2):67–??, Summer 2000. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/magnussonh2000:
67.html.

Normark:2000:EP

[150] Kurt Nørmark. Elucidative programming. Nordic Journal of Computing, 7(2):87–??, Summer 2000. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/normark:87.html.

Vestdam:2000:DTP

[151] Thomas Vestdam. Documentation threads — presentation of fragmented documentation. Nordic Journal of Computing, 7(2):106–??, Summer 2000. CODEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/vestdam2000:106.
                                        html.

Harsu:2000:IOO

[152] Maarit Harsu. Identifying objectoriented features from procedural software. Nordic Journal of Computing, 7(2):126–??, Summer 2000. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/harsu:126.html.

Bendix:2000:SSC

[153] Lars Bendix, Ulf Asklund, and Jonas Persson. Summary of the subworkshop on change management for open source software. Nordic Journal of Computing, 7(2):143–??, Summer 2000. CODEN NJ-COFR. ISSN 1236-6064.

Halldorsson:2000:GEF

[154] Mágnus M. Halldórsson. Guest Editor’s foreword. Nordic Journal of Computing, 7(3):149–??, Fall 2000. CODEN NJ-COFR. ISSN 1236-6064.

Pagh:2000:TWC

[155] Rasmus Pagh. A trade-off for worst-case efficient dictionaries. Nordic Journal of Computing, 7(3):151–??, Fall 2000. CODEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/pagh2000:151.html.

Berry:2000:RWT

[156] Anne Berry, Jean-Paul Bordat, and Pinar Heggernes. Recognizing weakly triangulated graphs by edge separability. Nordic Journal of Computing, 7(3):164–??, Fall 2000. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/berrybh2000:164.
html.

Berman:2000:AMW

[157] Piotr Berman. A d/2 approximation for maximum weight independent set in d-claw free graphs. Nordic Journal of Computing, 7(3):178–??, Fall 2000. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/berman:178.html.

Doddi:2000:AAC

[158] Srinivas Doddi, Madhav V. Marathe, S. S. Ravi, David S. Taylor, and Peter Widmayer. Approximation algorithms for clustering to minimize the sum of diameters. Nordic Journal of Computing, 7(3):185–??, Fall 2000. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/doddimrtw2000:
185.html.

Liberatore:2000:SJB

[159] Vincenzo Liberatore. Scheduling jobs before shut-down. Nordic Journal of Computing, 7(3):204–??, Fall 2000. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/liberatore2000:
204.html.

Agarwal:2000:PDT

[160] Pankaj K. Agarwal, Leonidas J. Guibas, Sariel Har-Peled, Alexander Rabinovitch, and Micha Sharir. Penetration depth of two convex polytopes in 3D. Nordic Journal of Computing, 7(3):227–??, Fall 2000. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/agarwalghrs2000:
                                        227.html.

Asano:2000:ORS

[161] Tetsuo Asano, Tomomi Matsui, and Takeshi Tokuyama. Optimal roundings of sequences and matrices. Nordic Journal of Computing, 7(3):241–??, Fall 2000. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/asanomt:241.html.

Smolka:2000:GEF

[162] Gert Smolka. Guest Editor’s foreword. Nordic Journal of Computing, 7(4):257– ??, Winter 2000. CODEN NJCOFR. ISSN 1236-6064.

Hofmann:2000:TSB

[163] Martin Hofmann. A type system for bounded space and functional in-place update. Nordic Journal of Computing, 7(4):258–??, Winter 2000. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/hofmann2000:258.
                                        html.

Mauborgne:2000:IUR

[164] Laurent Mauborgne. An incremental unique representation for regular trees. Nordic Journal of Computing, 7(4):290–??, Winter 2000. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/mauborgne2000:290.
html.

Pottier:2000:VCB

[165] François Pottier. A versatile constraintbased type inference system. Nordic Journal of Computing, 7(4):312–??, Winter 2000. CODEN NJCOFR. ISSN 1236-6064. URL http://www.cs.

helsinki.fi/njc/References/pottier2000:
312.html.

Russo:2000:FCS

[166] Claudio V. Russo. First-class structures for Standard ML. Nordic Journal of Computing, 7(4):348–??, Winter 2000. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/russo2000:348.
html.

Seidl:2000:CBI

[167] Helmut Seidl and Bernhard Steffen.

Constraint-based inter-procedural analysis of parallel programs. Nordic Journal of Computing, 7(4):375–??, Winter 2000. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/seidls:375.html.

Sere:2001:GEF

[168] Kaisa Sere and Wang Li. Guest Editors’ foreword. Nordic Journal of Computing, 8(1):1–??, Spring 2001. CODEN NJCOFR. ISSN 1236-6064.

Back:2001:CAS

[169] Ralph-Johan Back, Luigia Petre, and Ivan Porres. Continuous action systems as a model for hybrid systems. Nordic Journal of Computing, 8(1):2–??, Spring 2001. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/backpp2001:2.html.

Bove:2001:SGR

[170] Ana Bove. Simple general recursion in type theory. Nordic Journal of Computing, 8(1):22–??, Spring 2001. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/bove2001:22.html.

Hune:2001:GSC

[171] Thomas Hune, Kim G. Larsen, and Paul Pettersson. Guided synthesis of control programs using UPPAAL. Nordic Journal of Computing, 8(1):43–??, Spring 2001. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/hunelp2001:43.
html.

Jonsson:2001:IRS

[172] Bengt Jonsson, Tiziana Margaria, Gustaf Naeser, Jan Nyström, and Bernhard Steffen. Incremental requirement specification for evolving systems. Nordic Journal of Computing, 8(1):65–??, Spring 2001. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/jonssonmnns:65.
html.

Spelberg:2001:PRT

[173] R. F. Lutje Spelberg and W. J. Toetenel. Parametric real-time model checking using splitting trees. Nordic Journal of Computing, 8(1):88–??, Spring 2001. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/lutjespelbergt2001:
88.html.

Mork:2001:DIP

[174] Simon Mørk. Distributed implementation of a process-algebra based programming language for embedded systems. Nordic Journal of Computing, 8(1):121–??, Spring 2001. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/mork2001:121.html.

Ronkko:2001:LHA

[175] Mauno Rönkkö and Xuandong Li. Linear hybrid action systems. Nordic Journal of Computing, 8(1):159–??, Spring 2001. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/ronkko2001:159.
html.

Ottmann:2001:EEP

[176] Thomas Ottmann, Sven Schuierer, and Subbiah Soundaralakshmi. Enumerating extreme points in higher dimensions. Nordic Journal of Computing, 8(2):179–??, Summer 2001. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/ottmannss2001:179.
html.

Sere:2001:SVD

[177] Kaisa Sere and Marina Waldén. Structuring and verifying distributed algorithms. Nordic Journal of Computing, 8(2):193–??, Summer 2001. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/serew:193.html.

Gudmundsson:2001:AMM

[178] Joachim Gudmundsson, Christos Levcopoulos, and Giri Narasimhan. Approximating a minimum Manhattan network. Nordic Journal of Computing, 8(2):216–??, Summer 2001. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/gudmundssonln:216.
html.

Haveraaen:2001:GEF

[179] Magne Haveraaen and Olaf Owe. Guest Editors’ foreword. Nordic Journal of Computing, 8(3):277–??, Fall 2001. CO-DEN NJCOFR. ISSN 1236-6064.

Dosch:2001:CAF

[180] Walter Dosch and Sönke Magnussen. Computer aided fusion for algebraic program derivation. Nordic Journal of Computing, 8(3):279–??, Fall 2001. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/doschm2001:279.
html.

Lamo:2001:SPP

[181] Yngve Lamo and Michal Walicki. Specification of parameterized programs — persistency revisited. Nordic Journal of Computing, 8(3):298–??, Fall 2001. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/lamow2001:298.
html.

Petersson:2001:RSA

[182] Viktor Petersson and Sergei Vorobyov. A randomized subexponential algorithm for parity games. Nordic Journal of Computing, 8(3):324–??, Fall 2001. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/peterssonv2001:
324.html.

Fecher:2001:RTP

[183] Harald Fecher. A real-time process algebra with open intervals and maximal progress. Nordic Journal of Computing, 8(3):346–??, Fall 2001. CODEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.helsinki.
fi/njc/References/fecher2001:346.
html.

Uustalu:2001:RSC

[184] Tarmo Uustalu, Varmo Vene, and Alberto Pardo. Recursion schemes from comonads. Nordic Journal of Computing, 8(3):366–??, Fall 2001. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/uustaluvp2001:366.
html.

Gayo:2001:MDI

[185] Jose Emilio Labra Gayo, Juan Manuel Cueva Lovelle, Maria Cándida Luengo Díez, and Agustín Cernuda del Río. Modular development of interpreters from semantic building blocks. Nordic Journal of Computing, 8(3):391–407, Fall 2001. CODEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/labragayoclc2001:
391.html.

Bollobas:2001:TSS

[186] Béla Bollobás, Gautam Das, Dimitrios Gunopulos, and Heikki Mannila. Time-series similarity problems and well-separated geometric sets. Nordic Journal of Computing, 8(4):409–423, Winter 2001. CO-DEN NJCOFR. ISSN 1236-6064.

URL http://www.cs.helsinki.fi/
njc/References/bollobas2001:409.
html.

Ostergaard:2001:NAM

[187] Patric R. J. Östergård. A new algorithm for the maximum-weight clique problem. Nordic Journal of Computing, 8(4):424–436, Winter 2001. CO-DEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/ostergard2001:424.
                                        html.

Grobauer:2001:PEP

[188] Bernd Grobauer and Julia L. Lawall. Partial evaluation of pattern matching in strings, revisited. Nordic Journal of Computing, 8(4):437–462, Winter 2001. CODEN NJCOFR. ISSN 1236-6064.

                                        URL http://www.cs.helsinki.fi/
                                        njc/References/grobauer2001:437.
                                        html.

Boyar:2001:CRL

[189] Joan Boyar, Lene M. Favrholdt, Kim S. Larsen, and Morten N. Nielsen. The competitive ratio for on-line dual bin packing with restricted input sequences. Nordic Journal of Computing, 8(4):463–472, Winter 2001. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/boyar2001:463.
                                        html.

Burghardt:2001:MPS

[190] Jochen Burghardt. Maintaining partial sums in logarithmic time. Nordic Journal of Computing, 8(4):473–474, Winter 2001. CODEN NJCOFR. ISSN 1236-

                                        6064. URL http://www.cs.helsinki.
                                        fi/njc/References/burghardt2001:
                                        473.html.

Goldberg:2002:HCC

[191] M. Goldberg and M. Torgersen. How to circumvent Church numerals. Nordic Journal of Computing, 9(1):1–??, Spring 2002. CODEN NJCOFR. ISSN 1236-6064.

Narasimhan:2002:AAB

[192] G. Narasimhan and M. Smid. Approximation algorithms for the bottleneck stretch factor problem. Nordic Journal of Computing, 9(1):13–??, Spring 2002. CODEN NJCOFR. ISSN 1236-6064.

Hammar:2002:IER

[193] M. Hammar, B. J. Nilsson, and S. Schuierer. Improved exploration of rectilinear polygons. Nordic Journal of Computing, 9(1):32–??, Spring 2002. CODEN NJCOFR. ISSN 1236-6064.

Crochemore:2002:ASM

[194] M. Crochemore, C. Iliopoulos, C. Makris, W. Rytter, A. Tsakalidis, and K. Tsichlas. Approximate string matching with gaps. Nordic Journal of Computing, 9 (1):54–??, Spring 2002. CODEN NJ-COFR. ISSN 1236-6064.

Klipsch:2002:NBD

[195] K. N. Klipsch and D. S. Wise. Note:

Blood from Dahm’s turnip. Nordic Journal of Computing, 9(1):66–??, Spring 2002. CODEN NJCOFR. ISSN 1236-6064.

Aceto:2002:GEF

[196] L. Aceto and P. Panangaden. Guest Editors’ foreword. Nordic Journal of Computing, 9(2):69–??, Summer 2002. CO-DEN NJCOFR. ISSN 1236-6064. Selected papers of the Eighth International Workshop on Expressiveness in Concurrency (EXPRESS’01), August 20, 2001.

Amadio:2002:DCR

[197] R. M. Amadio and C. Meyssonnier. On decidability of the control reachability problem in the asynchronous picalculus. Nordic Journal of Computing, 9(2):70–??, Summer 2002. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Eighth International Workshop on Expressiveness in Concurrency (EXPRESS’01), August 20, 2001.

Bradfield:2002:IFM

[198] J. C. Bradfield and S. B. Fröschle. Independence-friendly modal logic and true concurrency. Nordic Journal of Computing, 9(2):102–??, Summer 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Eighth International Workshop on Expressiveness in Concurrency (EXPRESS’01), August 20, 2001.

Labroue:2002:ATA

[199] A. Labroue and Ph. Schnoebelen. An automata-theoretic approach to the reachability analysis of RPPS systems. Nordic Journal of Computing, 9(2): 118–??, Summer 2002. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Eighth International Workshop on Expressiveness in Concurrency (EXPRESS’01), August 20, 2001.

Nielsen:2002:TCC

[200] M. Nielsen, C. Palamidessi, and F. D. Valencia. Temporal concurrent constraint programming: Denotation, logic and applications. Nordic Journal of Computing, 9:145–??, Summer 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Eighth International Workshop on Expressiveness in Concurrency (EXPRESS’01), August 20, 2001.

Osterbye:2002:GEF

[201] K. Østerbye. Guest Editor’s foreword.

Nordic Journal of Computing, 9(3):189– ??, Fall 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Tenth Nordic Workshop on Programming Environment Research (NW-PER’2002), August 18–20, 2002.

Ernst:2002:SDM

[202] E. Ernst. Safe dynamic multiple inheritance. Nordic Journal of Computing, 9(3):191–??, Fall 2002. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Tenth Nordic Workshop on Programming Environment Research (NW-PER’2002), August 18–20, 2002.

Vestdam:2002:EPT

[203] T. Vestdam. Elucidative program tutorials. Nordic Journal of Computing, 9(3):209–??, Fall 2002. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Tenth Nordic Workshop on Programming Environment Research (NW-PER’2002), August 18–20, 2002.

Arisholm:2002:WBS

[204] E. Arisholm, D. I. K. Sjøberg, G. J.

Carelius, and Y. Lindsjørn. A Webbased support environment for software engineering experiments. Nordic Journal of Computing, 9(3):231–??, Fall 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Tenth Nordic Workshop on Programming Environment Research (NWPER’2002), August 18–20, 2002.

Hammouda:2002:PBJ

[205] I. Hammouda and K. Koskimies. A pattern-based J2EE application development environment. Nordic Journal of Computing, 9(3):248–??, Fall 2002. CO-DEN NJCOFR. ISSN 1236-6064. Selected papers of the Tenth Nordic Workshop on Programming Environment Research (NWPER’2002), August 18–20, 2002.

Bendix:2002:SSE

[206] L. Bendix and G. Hedin. Summary of the Subworkshop on Extreme Programming. Nordic Journal of Computing, 9:261–??, Fall 2002. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Tenth Nordic Workshop on Programming Environment Research (NW-PER’2002), August 18–20, 2002.

Haveraaen:2002:GEF

[207] M. Haveraaen and M. R. Hansen. Guest Editors’ foreword. Nordic Journal of Computing, 9(4):267–??, ???? 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Thirteenth Nordic Workshop on Programming Theory (NWPT’01), October 10–12, 2001.

Amnell:2002:CST

[208] T. Amnell, E. Fersman, P. Pettersson, H. Sun, and W. Yi. Code synthesis for timed automata. Nordic Journal of Computing, 9(4):269–??, ???? 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Thirteenth Nordic Workshop on Programming Theory (NWPT’01), October 10–12, 2001.

Hoenicke:2002:COD

[209] J. Hoenicke and E.-R. Olderog. CSP-OZ-DC: A combination of specification techniques for processes, data and time. Nordic Journal of Computing, 9(4):301– ??, ???? 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Thirteenth Nordic Workshop on Programming Theory (NWPT’01), October 10– 12, 2001.

Nielson:2002:SSA

[210] F. Nielson, H. Seidl, and H. Riis Nielson. A succinct solver for ALFP. Nordic Journal of Computing, 9(4):335–??, ???? 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Thirteenth Nordic Workshop on Programming Theory (NWPT’01), October 10–12, 2001.

Wagner:2002:ASS

[211] E. G. Wagner. Algebraic specifications:

some old history and new thoughts. Nordic Journal of Computing, 9(4):373– ??, ???? 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Thirteenth Nordic Workshop on Programming Theory (NWPT’01), October 10– 12, 2001.

Wehrheim:2002:RSB

[212] H. Wehrheim. Relating state-based and behaviour-oriented subtyping. Nordic Journal of Computing, 9(4):405–??, ???? 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Thirteenth Nordic Workshop on Programming Theory (NWPT’01), October 10–12, 2001.

Anonymous:2002:AIV

[213] Anonymous. Author index volume 9 (2002). Nordic Journal of Computing, 9(4):436–??, ???? 2002. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Thirteenth Nordic Workshop on Programming Theory (NWPT’01), October 10–12, 2001.

Holub:2003:GEF

[214] J. Holub. Guest Editor’s foreword. Nordic Journal of Computing, 10(1): 1, Spring 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23–24, 2002.

Baba:2003:NRA

[215] K. Baba, A. Shinohara, M. Takeda, S. Inenaga, and S. Arikawa. A note on randomized algorithm for string matching with mismatches. Nordic Journal of Computing, 10(1):2–??, Spring 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23–24, 2002.

Cinque:2003:WOP

[216] L. Cinque, S. De Agostino, and F. Liberati. A work-optimal parallel implementation of lossless image compression by block matching. Nordic Journal of Computing, 10(1):13–??, Spring 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23–24, 2002.

Franek:2003:NCR

[217] F. Franek, W. F. Smyth, and X. Xiao. A note on Crochemore’s repetitions algorithm — A fast space-efficient approach. Nordic Journal of Computing, 10(1):21–??, Spring 2003. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23–24, 2002.

Hyyro:2003:BVA

[218] H. Hyyrö. A bit-vector algorithm for computing Levenshtein and Damerau edit distances. Nordic Journal of Computing, 10(1):29–??, Spring 2003. CO-DEN NJCOFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23–24, 2002.

Iliopoulos:2003:SRD

[219] C. S. Iliopoulos, M. Mohamed, L. Mouchard, K. G. Perdikuri, W. F. Smyth, and A. K. Tsakalidis. String regularities with don’t cares. Nordic Journal of Computing, 10 (1):40–??, Spring 2003. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23–24, 2002.

Inenaga:2003:BCS

[220] S. Inenaga. Bidirectional construction of suffix trees. Nordic Journal of Computing, 10(1):52–??, Spring 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23–24, 2002.

Nestmann:2003:GEF

[221] U. Nestmann and P. Panangaden. Guest Editors’ foreword. Nordic Journal of Computing, 10(2):69–??, Summer 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Ninth International Workshop on Expressiveness in Concurrency (EXPRESS’02), August 19, 2002.

Carbone:2003:EPP

[222] M. Carbone and S. Maffeis. On the expressive power of polyadic synchronisation in pi-calculus. Nordic Journal of Computing, 10(2):70–??, Summer 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Ninth International Workshop on Expressiveness in Concurrency (EXPRESS’02), August 19, 2002.

Ouaknine:2003:TCC

[223] J. Ouaknine and J. Worrell. Timed CSP = closed timed epsilon-automata. Nordic Journal of Computing, 10(2): 99–??, Summer 2003. CODEN NJ-COFR. ISSN 1236-6064. Selected papers of the Ninth International Workshop on Expressiveness in Concurrency (EXPRESS’02), August 19, 2002.

Fernandez:2003:INM

[224] M. Fernández and L. Khalil. Interaction nets with McCarthy’s amb: Properties and applications. Nordic Journal of Computing, 10(2):134–??, Summer 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Ninth International Workshop on Expressiveness in Concurrency (EXPRESS’02), August 19, 2002.

Sassone:2003:DBC

[225] V. Sassone and P. Sobocinski. Deriving bisimulation congruences using 2categories. Nordic Journal of Computing, 10(2):163–??, Summer 2003. CO-DEN NJCOFR. ISSN 1236-6064. Selected papers of the Ninth International Workshop on Expressiveness in Concurrency (EXPRESS’02), August 19, 2002.

Lemstrom:2003:TIP

[226] K. Lemström and J. Tarhio. Transposition invariant pattern matching for multi-track strings. Nordic Journal of Computing, 10(3):185–??, Fall 2003. CODEN NJCOFR. ISSN 1236-6064.

Fiala:2003:GCC

[227] J. Fiala, P. Heggernes, P. Kristiansen, and J. A. Telle. Generalized H-coloring and H-covering of trees. Nordic Journal of Computing, 10(3):206–??, Fall 2003. CODEN NJCOFR. ISSN 1236-6064.

Segal:2003:POF

[228] M. Segal. Placing an obnoxious facility in geometric networks. Nordic Journal of Computing, 10(3):225–??, Fall 2003. CODEN NJCOFR. ISSN 1236-6064.

Katajainen:2003:NPA

[229] J. Katajainen and F. Vitale. Navigation piles with applications to sorting, priority queues, and priority deques. Nordic Journal of Computing, 10(3):238–??, Fall 2003. CODEN NJ-COFR. ISSN 1236-6064.

Haveraaen:2003:GEF

[230] M. Haveraaen and J. Vain. Guest Editors’ foreword. Nordic Journal of Computing, 10(4):263–??, Winter 2003. CO-DEN NJCOFR. ISSN 1236-6064.

Benke:2003:UGP

[231] M. Benke, P. Dybjer, and P. Jansson. Universes for generic programs and proofs in dependent type theory. Nordic Journal of Computing, 10(4): 265–??, Winter 2003. CODEN NJ-COFR. ISSN 1236-6064.

Ghani:2003:RC

[232] N. Ghani and C. Lüth. Rewriting via coinserters. Nordic Journal of Computing, 10(4):290–??, Winter 2003. CODEN NJCOFR. ISSN 1236-6064.

Johnsen:2003:ART

[233] E. B. Johnsen and C. Lüth. Abstracting refinements for transformation. Nordic Journal of Computing, 10(4):313–??, Winter 2003. CODEN NJCOFR. ISSN 1236-6064.

Pilegaard:2003:AAA

[234] H. Pilegaard, M. R. Hansen, and R. Sharp. An approach to analyzing availability properties of security protocols. Nordic Journal of Computing, 10 (4):337–??, Winter 2003. CODEN NJ-COFR. ISSN 1236-6064.

Anonymous:2003:AIV

[235] Anonymous. Author index volume 10 (2003). Nordic Journal of Computing, 10(4):374–??, Winter 2003. CODEN NJ-COFR. ISSN 1236-6064.

Gairing:2004:SSA

[236] M. Gairing, R. M. Geist, S. T. Hedetniemi, and P. Kristiansen. A selfstabilizing algorithm for maximal 2packing. Nordic Journal of Computing, 11(1):1–??, Spring 2004. CODEN NJ-COFR. ISSN 1236-6064.

Sivertsen:2004:UVU

[237] T. Sivertsen. Undefinedness vs. underspecification in HALDEN ASL. Nordic Journal of Computing, 11(1):12–??, Spring 2004. CODEN NJCOFR. ISSN 1236-6064.

Tronicek:2004:EDA

[238] Z. Tronícek. Episode directed acyclic subsequence graph. Nordic Journal of Computing, 11(1):35–??, Spring 2004. CODEN NJCOFR. ISSN 1236-6064.

Berlea:2004:BQD

[239] A. Berlea and H. Seidl. Binary queries for document trees. Nordic Journal of Computing, 11(1):41–??, Spring 2004. CODEN NJCOFR. ISSN 1236-6064.

Sere:2004:GEF

[240] K. Sere and M. Waldén. Guest Editors’ foreword. Nordic Journal of Computing, 11(2):73–??, Summer 2004. CODEN NJ-COFR. ISSN 1236-6064.

Bostrom:2004:ICS

[241] P. Boström and M. Waldén. Implementation of control systems using B action systems: A case study. Nordic Journal of Computing, 11(2):75–??, Summer 2004. CODEN NJCOFR. ISSN 1236-6064.

Celiku:2004:CBA

[242] O. Celiku and A. McIver. Cost-based analysis of probabilistic programs mechanised in HOL. Nordic Journal of Computing, 11(2):102–??, Summer 2004. CODEN NJCOFR. ISSN 1236-6064.

Fersman:2004:GAS

[243] E. Fersman and W. Yi. A generic approach to schedulability analysis of realtime tasks. Nordic Journal of Computing, 11(2):129–??, Summer 2004. CO-DEN NJCOFR. ISSN 1236-6064.

Latvala:2004:IFV

[244] T. Latvala and H. Tauriainen. Improved on-the-fly verification with testers. Nordic Journal of Computing, 11(2): 148–??, Summer 2004. CODEN NJ-COFR. ISSN 1236-6064.

Ruksenas:2004:RED

[245] R. Ruksenas. A rigourous environment for development of concurrent systems. Nordic Journal of Computing, 11(2): 165–??, Summer 2004. CODEN NJ-COFR. ISSN 1236-6064.

Schneider:2004:CIK

[246] G. Schneider. Computing invariance kernels of polygonal hybrid systems. Nordic Journal of Computing, 11(2): 194–??, Summer 2004. CODEN NJ-COFR. ISSN 1236-6064.

Lilius:2004:GEF

[247] J. Lilius and K. Østerbye. Guest Editors’ foreword: Programming and software development environment research in the Nordic countries. Nordic Journal of Computing, 11(3):211–??, Fall 2004. CODEN NJCOFR. ISSN 1236-6064.

Nilsson:2004:IJC

[248] A. Nilsson, A. Ive, T. Ekman, and G. Hedin. Implementing Java compilers using ReRAGs. Nordic Journal of Computing, 11(3):213–??, Fall 2004. CODEN NJCOFR. ISSN 1236-6064.

Selonen:2004:TUP

[249] P. Selonen, M. Siikarla, K. Koskimies, and T. Mikkonen. Towards the unification of patterns and profiles in UML. Nordic Journal of Computing, 11(3): 235–??, Fall 2004. CODEN NJCOFR. ISSN 1236-6064.

Staron:2004:CSP

[250] M. Staron, L. Kuzniarz, and L. Wallin. Case study on a process of industrial MDA realization – determinants of effectiveness. Nordic Journal of Computing, 11(3):254–??, Fall 2004. CODEN NJCOFR. ISSN 1236-6064.

Storrle:2004:SNU

[251] H. Störrle. Structured nodes in UML 2.0 activities. Nordic Journal of Computing, 11(3):279–??, Fall 2004. CODEN NJCOFR. ISSN 1236-6064.

Vestdam:2004:MPU

[252] T. Vestdam and K. Nørmark. Maintaining program understanding – issues, tools, and future directions. Nordic Journal of Computing, 11(3):303–??, Fall 2004. CODEN NJCOFR. ISSN 1236-6064.

Sutinen:2004:ASM

[253] E. Sutinen and J. Tarhio. Approximate string matching with ordered qgrams. Nordic Journal of Computing, 11 (4):321–??, Winter 2004. CODEN NJ-COFR. ISSN 1236-6064.

Bereg:2004:DAA

[254] S. Bereg and M. Segal. Dynamic algorithms for approximating interdistances. Nordic Journal of Computing, 11(4): 344–??, Winter 2004. CODEN NJ-COFR. ISSN 1236-6064.

Navarro:2004:ARE

[255] G. Navarro. Approximate regular expression searching with arbitrary integer weights. Nordic Journal of Computing, 11(4):356–??, Winter 2004. CODEN NJ-COFR. ISSN 1236-6064.

Bodlaender:2004:SEC

[256] H. L. Bodlaender and J. A. Telle. Spaceefficient construction variants of dynamic programming. Nordic Journal of Computing, 11(4):374–??, Winter 2004. CODEN NJCOFR. ISSN 1236-6064.

Anonymous:2004:AIV

[257] Anonymous. Author index volume 11 (2004). Nordic Journal of Computing, 11(4):386–??, Winter 2004. CODEN NJ-COFR. ISSN 1236-6064.

Krizanc:2005:RMR

[258] Danny Krizanc, Pat Morin, and Michiel Smid. Range mode and range median queries on lists and trees. Nordic Journal of Computing, 12(1):1–17, Spring 2005. CODEN NJCOFR. ISSN 1236-6064.

Fokkink:2005:NKS

[259] Wan Fokkink, Jaap-Henk Hoepman, and Jun Pang. A note on K-state selfstabilization in a ring with K = N. Nordic Journal of Computing, 12(1):18– 26, Spring 2005. CODEN NJCOFR. ISSN 1236-6064.

Elmasry:2005:DJ

[260] Amr Elmasry. Deterministic jumplists. Nordic Journal of Computing, 12(1):27– 39, Spring 2005. CODEN NJCOFR. ISSN 1236-6064.

Makinen:2005:SSA

[261] Veli Mäkinen and Gonzalo Navarro. Succinct suffix arrays based on run-length encoding. Nordic Journal of Computing, 12(1):40–66, Spring 2005. CODEN NJ-COFR. ISSN 1236-6064.

Pettersson:2005:GEF

[262] Paul Pettersson and Wang Yi. Guest Editors’ foreword. Nordic Journal of Computing, 12(2):67–??, Summer 2005. CODEN NJCOFR. ISSN 1236-6064.

Ernits:2005:MAS

[263] Juhan Ernits. Memory arbiter synthesis and verification for a radar memory interface card. Nordic Journal of Computing, 12(2):68–88, Summer 2005. CO-DEN NJCOFR. ISSN 1236-6064.

Kristiansen:2005:PLC

[264] Lars Kristiansen and Paul J. Voda. Programming languages capturing complexity classes. Nordic Journal of Computing, 12(2):89–115, Summer 2005. CO-DEN NJCOFR. ISSN 1236-6064.

Kyas:2005:CTL

[265] Marcel Kyas, Frank S. de Boer, and Willem-Paul de Roever. A compositional trace logic for behavioural interface specifications. Nordic Journal of Computing, 12(2):116–132, Summer 2005. CODEN NJCOFR. ISSN 1236-6064.

Nestra:2005:TC

[266] Härmel Nestra. Transfinite corecursion.

Nordic Journal of Computing, 12(2): 133–156, Summer 2005. CODEN NJ-COFR. ISSN 1236-6064.

Runde:2005:RUI

[267] Ragnhild Kobro Runde, Øystein Haugen, and Ketil Stølen. Refining UML interactions with underspecification and nondeterminism. Nordic Journal of Computing, 12(2):157–188, Summer 2005. CODEN NJCOFR. ISSN 1236-6064.

Veldhuizen:2005:LEP

[268] Todd L. Veldhuizen. Language embeddings that preserve staging and safety. Nordic Journal of Computing, 12(2): 189–198, Summer 2005. CODEN NJ-COFR. ISSN 1236-6064.

Lilius:2005:GEF

[269] Johan Lilius, Ricardo J. Machado, Dragos Truscan, João M. Fernandes, and Ivan Porres. Guest Editors’ foreword. Nordic Journal of Computing, 12(3): 198–199, Fall 2005. CODEN NJCOFR. ISSN 1236-6064.

Al-Azzoni:2005:MVC

[270] Issam Al-Azzoni, Douglas G. Down, and Ridha Khedri. Modeling and verification of cryptographic protocols using coloured Petri nets and design/CPN. Nordic Journal of Computing, 12(3): 200–228, Fall 2005. CODEN NJCOFR. ISSN 1236-6064.

Barros:2005:CSC

[271] João Paulo Barros and Jens Bæk Jørgensen. A case study on coloured Petri nets in object-oriented analysis and design. Nordic Journal of Computing, 12(3):229–250, Fall 2005. CODEN NJCOFR. ISSN 1236-6064.

Billington:2005:VCS

[272] Jonathan Billington, Guy Edward Gallasch, and Laure Petrucci. Verification of the class of stop-and-wait protocols modelled by coloured Petri nets. Nordic Journal of Computing, 12(3):251–274, Fall 2005. CODEN NJCOFR. ISSN 1236-6064.

Petrank:2005:HCC

[273] Erez Petrank and Dror Rawitz. The hardness of cache conscious data placement. Nordic Journal of Computing, 12 (3):275–307, Fall 2005. CODEN NJ-COFR. ISSN 1236-6064.

Prieto:2005:RIS

[274] Elena Prieto and Christian Sloper. Reducing to independent set structure – the case of k-internal spanning tree. Nordic Journal of Computing, 12(3): 308–318, Fall 2005. CODEN NJCOFR. ISSN 1236-6064.

Koskimies:2005:GEF

[275] Kai Koskimies and Ludwik Kuźniarz. Guest Editors’ foreword. Nordic Journal of Computing, 12(4):319–320, Winter 2005. CODEN NJCOFR. ISSN 1236-6064.

Alanen:2005:CMF

[276] Marcus Alanen, Torbjörn Lundkvist, and Ivan Porres. Comparison of modeling frameworks for software engineering. Nordic Journal of Computing, 12 (4):321–342, Winter 2005. CODEN NJ-COFR. ISSN 1236-6064.

Dobrzanski:2005:PRE

[277] Łukasz Dobrzański and Ludwik Kuźniarz.

Practical refactoring of executable UML models. Nordic Journal of Computing, 12(4):343–360, Winter 2005. CODEN NJCOFR. ISSN 1236-6064.

Lilius:2005:AEE

[278] Johan Lilius, Tomas Lillqvist, Torbjörn Lundkvist, Ian Oliver, Ivan Porres, Kim Sandström, Glen Sveholm, and Asim Pervez Zaka. An architecture exploration environment for system on chip design. Nordic Journal of Computing, 12(4):361–378, Winter 2005. CO-DEN NJCOFR. ISSN 1236-6064.

Wenzel:2005:ADI

[279] Sven Wenzel. Automatic detection of incomplete instances of structural patterns in UML class diagrams. Nordic Journal of Computing, 12(4):361–378, Winter 2005. CODEN NJCOFR. ISSN 1236-6064.

Jones:2006:GEF

[280] Neil D. Jones. Guest Editor’s foreword.

Nordic Journal of Computing, 13(1-2): 1–??, Summer 2006. CODEN NJCOFR. ISSN 1236-6064.

Bortin:2006:SFD

[281] Maksym Bortin, Einar Broch Johnsen, and Christoph Lüth. Structured formal development in Isabelle. Nordic Journal of Computing, 13(1-2):2–21, Summer 2006. CODEN NJCOFR. ISSN 1236-6064.

Bruckner:2006:OSF

[282] Ingo Brückner, Björn Metzler, and Heike Wehrheim. Optimizing slicing of formal specifications by deductive verification. Nordic Journal of Computing, 13 (1-2):22–45, Summer 2006. CODEN NJ-COFR. ISSN 1236-6064.

Crolard:2006:EPL

[283] Tristan Crolard, Samuel Lacas, and Pierre Valarcher. On the expressive power of the loop language. Nordic Journal of Computing, 13(1-2):46–57, Summer 2006. CODEN NJCOFR. ISSN 1236-6064.

Damgaard:2006:ABB

[284] Troels C. Damgaard and Lars Birkedal. Axiomatizing binding bigraphs. Nordic Journal of Computing, 13(1-2):58–77, Summer 2006. CODEN NJCOFR. ISSN 1236-6064.

Nesi:2006:DTF

[285] Monica Nesi and Giustina Nocera. Deriving the type flaw attacks in the Otway–Rees protocol by rewriting. Nordic Journal of Computing, 13(1-2): 78–97, Summer 2006. CODEN NJ-COFR. ISSN 1236-6064.

Nielsen:2006:SAB

[286] Christoffer Rosenkilde Nielsen and Hanne Riis Nielson. Static analysis for blinding. Nordic Journal of Computing, 13(1-2):98–116, Summer 2006. CODEN NJCOFR. ISSN 1236-6064.

Stovring:2006:HOB

[287] Kristian Støvring. Higher-order beta matching with solutions in long beta-eta normal form. Nordic Journal of Computing, 13(1-2):117–126, Summer 2006. CODEN NJCOFR. ISSN 1236-6064.

Tsiopoulos:2006:FDN

[288] Leonidas Tsiopoulos and Marina Waldén. Formal development of NoC systems in B. Nordic Journal of Computing, 13(1-2):127–145, Summer 2006. CODEN NJ-COFR. ISSN 1236-6064.

Prinz:2006:GEF

[289] Andreas Prinz and Merete Skjelten Tveit. Guest Editor’s foreword. Nordic Journal of Computing, 13(3):147–148, Fall 2006. CODEN NJCOFR. ISSN 1236-6064.

Byckling:2006:ICO

[290] Pauli Byckling, Petri Gerdt, Ludwik Kuzniarz, and Jorma Sajaniemi. Increasing comprehensibility of object models: Making the roles of attributes explicit in UML diagrams. Nordic Journal of Computing, 13(3):149–161, Fall 2006. CODEN NJCOFR. ISSN 1236-6064.

Koskinen:2006:PBA

[291] Johannes Koskinen, Anna Ruokonen, and Tarja Systä. A pattern-based approach to generate code from API usage scenarios. Nordic Journal of Computing, 13(3):162–179, Fall 2006. CODEN NJ-COFR. ISSN 1236-6064.

Siikarla:2006:TUM

[292] Mika Siikarla, Jari Peltonen, and Johannes Koskinen. Towards unambiguous model fragments. Nordic Journal of Computing, 13(3):180–195, Fall 2006. CODEN NJCOFR. ISSN 1236-6064.

Santos:2006:MDA

[293] André L. Santos, Kai Koskimies, and Antónia Lopes. A model-driven approach to variability management in product-line engineering. Nordic Journal of Computing, 13(3):196–213, Fall 2006. CODEN NJCOFR. ISSN 1236-

Gebremedhin:2006:PMD

[294] Assefaw Hadish Gebremedhin, Mohamed Essaïdi, Isabelle Guérin Lassous, Jens Gustedt, and Jan Arne Telle. PRO: A model for the design and analysis of efficient and scalable parallel algorithms. Nordic Journal of Computing, 13(4):215–239, Winter 2006. CODEN NJCOFR. ISSN 1236-6064.

Berlea:2006:OER

[295] Alexandru Berlea. Online evaluation of regular tree queries. Nordic Journal of Computing, 13(4):240–265, Winter 2006. CODEN NJCOFR. ISSN 1236-6064.

Poddar:2006:VGB

[296] Rajiv Kumar Poddar and Purandar Bhaduri. Verification of giotto based embedded control systems. Nordic Journal of Computing, 13(4):266–293, Winter 2006. CODEN NJCOFR. ISSN 1236-6064.

Gupta:2006:RAQ

[297] Prosenjit Gupta. Range-aggregate query problems involving geometric aggregation operations. Nordic Journal of Computing, 13(4):294–308, Winter 2006. CODEN NJCOFR. ISSN 1236-6064.

Subramani:2006:CPD

[298] K. Subramani and John Argentieri. Chain programming over difference constraints. Nordic Journal of Computing, 13(4):309–327, Winter 2006. CODEN NJCOFR. ISSN 1236-6064.

Lemire:2006:SMM

[299] Daniel Lemire. Streaming maximumminimum filter using no more than three comparisons per element. Nordic Journal of Computing, 13(4):328–339, Winter 2006. CODEN NJCOFR. ISSN 1236-6064.

Schellekens:2006:SDT

[300] Michel Schellekens, Rachit Agarwal, Emanuel Popovici, and Ka Lok Man. A simplified derivation of timing complexity lower bounds for sorting by comparisons. Nordic Journal of Computing, 13 (4):340–346, Winter 2006. CODEN NJ-COFR. ISSN 1236-6064.

Abdulla:2007:UFR

[301] Parosh Aziz Abdulla, Johann Deneux, Pritha Mahata, and Aletta Nylén. Using forward reachability analysis for verification of timed Petri nets. Nordic Journal of Computing, 14(1-2):1–42, Summer 2007. CODEN NJCOFR. ISSN 1236-6064.

Lam:2007:FRA

[302] Vitus S. W. Lam. A formalism for reasoning about UML activity diagrams. Nordic Journal of Computing, 14(1-2): 43–64, Summer 2007. CODEN NJ-COFR. ISSN 1236-6064.

Man:2007:FSA

[303] Ka Lok Man. Formal specification and analysis of hardware systems in Timed Chi. Nordic Journal of Computing, 14 (1-2):65–86, Summer 2007. CODEN NJ-COFR. ISSN 1236-6064.

Heggernes:2007:LTC

[304] Pinar Heggernes and Dieter Kratsch.

Linear-time certifying recognition algorithms and forbidden induced subgraphs. Nordic Journal of Computing, 14(1-2):87–108, Summer 2007. CODEN NJCOFR. ISSN 1236-6064.

Bremler-Barr:2007:PLT

[305] Anat Bremler-Barr and Leah Epstein. Path layout on tree networks: Bounds in different label switching models. Nordic Journal of Computing, 14(1-2):109–125, Summer 2007. CODEN NJCOFR. ISSN 1236-6064.

Mahajan:2007:BSC

[306] Meena Mahajan, Raghavan Rama, and S. Vijayakumar. Block sorting: A characterization and some heuristics. Nordic Journal of Computing, 14(1-2):126–150, Summer 2007. CODEN NJCOFR. ISSN 1236-6064.

Cieliebak:2008:CVE

[307] Mark Cieliebak, Stephan Eidenbenz, Aris T. Pagourtzis, and Konrad Schlude. On the complexity of variations of equal sum subsets. Nordic Journal of Computing, 14(3):151–172, Fall 2008. CODEN NJCOFR. ISSN 1236-6064.

Xu:2008:FST

[308] Qiwen Xu and Naijun Zhan. Formalising scheduling theories in duration calculus. Nordic Journal of Computing, 14(3):173–201, Fall 2008. CODEN NJ-COFR. ISSN 1236-6064.

Kranakis:2008:TMU

[309] Evangelos Kranakis, Danny Krizanc, and Sunil Shende. Tracking mobile users in cellular networks using timing information. Nordic Journal of Computing, 14(3):202–215, Fall 2008. CODEN NJ-COFR. ISSN 1236-6064.

Boichut:2008:ABT

[310] Y. Boichut, P.-C. Héam, and O. Kouchnarenko. Approximation-based tree regular modelchecking. Nordic Journal of Computing, 14(3):216–241, Fall 2008. CODEN NJ-COFR. ISSN 1236-6064.

Hvannberg:2008:GEF

[311] Ebba Hvannberg. Guest Editor’s foreword. Nordic Journal of Computing, 14 (4):243–244, Winter 2008. CODEN NJ-COFR. ISSN 1236-6064.

Carrez:2008:SES

[312] Cyril Carrez, Lotte Johansen, Pawel Cieslak, and Stefan Hänsgen. Service engineering with the SIMS design and validation tools. Nordic Journal of Computing, 14(4):245–263, Winter 2008. CO-DEN NJCOFR. ISSN 1236-6064.

Lahtinen:2008:MBA

[313] Samuel Lahtinen and Kai Koskimies.

A model-based approach to reflective command interfaces. Nordic Journal of Computing, 14(4):264–281, Winter 2008. CODEN NJCOFR. ISSN 1236-6064.

Lind:2008:ERT

[314] Kenneth Lind and Rogardt Heldal. Estimation of real-time software component size. Nordic Journal of Computing, 14 (4):282–300, Winter 2008. CODEN NJ-COFR. ISSN 1236-6064.

Mohagheghi:2008:MSP

[315] Parastoo Mohagheghi, Vegard Dehlen, and Tor Neple. A metamodel and supporting process and tool for specifying quality models in model-based software development. Nordic Journal of Computing, 14(4):301–320, Winter 2008. CODEN NJCOFR. ISSN 1236-6064.

Raatikainen:2008:FME

[316] Mikko Raatikainen, Varvana Myllärniemi, and Tomi Männistö. Featback: Method for enhancing management of agile development. Nordic Journal of Computing, 14(4):321–337, Winter 2008. CO-DEN NJCOFR. ISSN 1236-6064.

Raiha:2008:PBG

[317] Outi Räihä, Kai Koskimies, Erkki Mäkinen, and Tarja Systä. Patternbased genetic model refinements in MDA. Nordic Journal of Computing, 14 (4):338–355, Winter 2008. CODEN NJ-COFR. ISSN 1236-6064.

Peltonen:2013:GEF

[318] Jari Peltonen. Guest Editor’s foreword. Nordic Journal of Computing, 15(1):1–2, Spring 2013. CODEN NJCOFR. ISSN 1236-6064.

Harsu:2013:CCG

[319] Maarit Harsu, Tibor Bakota, István Siket, Kai Koskimies, and Tarja Systä. Code clones: Good, bad, or ugly? Nordic Journal of Computing, 15(1):3– 17, Spring 2013. CODEN NJCOFR. ISSN 1236-6064.

Herczeg:2013:TFF

[320] Zoltán Herczeg, Gábor Lóki, Tamás Szirbucz, and Ákos Kiss. Validating JavaScript guidelines across multiple Web browsers. Nordic Journal of Computing, 15(1):18–31, Summer 2013. CO-DEN NJCOFR. ISSN 1236-6064.

Sakkinen:2013:TFF

[321] Markku Sakkinen, Philippe Lahire, and Ciprian-Bogdan Chirilă. Towards fullyfledged reverse inheritance in Eiffel. Nordic Journal of Computing, 15(1):32– 52, Spring 2013. CODEN NJCOFR. ISSN 1236-6064.

Tiusanen:2013:GKA

[322] Mikko Tiusanen and Antti Valmari.

Good to know about the efficiency of state space methods. Nordic Journal of Computing, 15(1):53–74, Spring 2013. CODEN NJCOFR. ISSN 1236-6064.